### v1

# Absolute concentration robustness: algebra and geometry

Abstract.
Motivated by the question of how biological systems maintain homeostasis in changing environments, Shinar and Feinberg introduced in 2010 the concept of absolute concentration robustness (ACR). A biochemical system exhibits ACR in some species if the steady-state value of that species does not depend on initial conditions. Thus, a system with ACR can maintain a constant level of one species even as the environment changes. Despite a great deal of interest in ACR in recent years, the following basic question remains open: How can we determine quickly whether a given biochemical system has ACR? Although various approaches to this problem have been proposed, we show that they are incomplete. Accordingly, we present new methods for deciding ACR, which harness computational algebra. We illustrate our results on several biochemical signaling networks.
MSC Codes: 37N25, 92E20, 12D10, 37C25, 65H14, 14Q20
1. Introduction
How do cells maintain function in fluctuating environments? In signaling transduction pathways, such fluctuations might be in the abundances of signaling proteins or may arise from crosstalk between proteins. Nevertheless, robustness of an output signal has been observed experimentally. In E. coli, for instance, such robustness is found in the EnvZ-OmpR osmoregularity signaling pathway [5, 69] and the glyoxylate bypass mechanism [47, 70], often over several orders of magnitude in the abundances of individual signaling proteins.
Mathematically, this phenomenon was first rigorously studied by Shinar and Feinberg who introduced the concept of absolute concentration robustness (ACR) [68]. A biochemical system exhibits ACR in some species $X_{i}$ if at every positive steady state, regardless of initial conditions, the concentration of $X_{i}$ is the same. Thus, a system with ACR maintains species $X_{i}$ at a constant level, even when the environment changes.
Most prior results regarding ACR are sufficient conditions for ACR. The first such criterion, due to Shinar and Feinberg, can be verified easily (as implemented in [46]) from the structure of the network of biochemical interactions [68]. Notably, this sufficient condition for ACR is also necessary for some classes of networks [52]. Other criteria for ACR were given by Neigenfind et al., who harnessed methods based on elementary modes [58, 59], and by Tonello and Johnston, who used the theory of network translation [35, 75].
Other studies have focused on developing and applying results on ACR to understand the biochemical mechanisms underlying robustness, for example, in mechanisms with dimerization and bifunctional proteins [19, 20]. This topic of biological robustness has generated interest in ACR from control theory [11, 43] and also inspired Craciun and Joshi to refine the definition of ACR to better capture the types of robustness relevant to applications [37, 38]. Additionally, the property of ACR has been studied in randomly generated networks [39] and in models that are stochastic, rather than deterministic [1, 2, 25].
ACR has also been analyzed using methods from computational algebra [41, 42, 63]. Our work builds on those works to relate ACR to various algebraic objects (ideals and varieties) associated to a (biochemical or mass-action) system. We also distinguish between ACR of a given system and ACR for a family of systems arising from a given reaction network. More generally, we explore the underlying algebraic and geometric structure of ACR.
We also consider the problem of deciding whether a network has ACR. We first show that for two classes of networks common in applications – those admitting a steady-state parameterization or having no conservation laws – this decision problem is easy (Section 3). Subsequently, we show by counterexample that various approaches (some in the literature) to this decision problem are incomplete (Section 4.4). Finally, we develop new procedures, using computer algebra, to tackle this problem. Specifically, we use a new ideal, which we call the positive-restriction ideal (Section 4), numerical algebraic geometry (Section 5), and specializations of Gröbner bases (Section 6).
We are not the first to harness numerical algebraic geometry [30, 57] and specializations of Gröbner bases [22] to prove results about biochemical systems. Indeed, a great deal of research has been conducted recently using algebraic techniques to analyze steady states of mass-action systems. Many such investigations concern when steady states are defined by binomial equations  [17, 35, 53]. Other studies have focused on computing steady-state invariants [42], constructing and harnessing steady-state parameterizations [12, 22, 28, 36, 62, 64], computing the maximum number of steady states [30, 61], and characterizing bistability and oscillations [13, 14, 76].
The outline of our work is as follows. We introduce mass-action systems and ACR in Section 2. We then look at the relationship between ACR and multistationarity, steady-state parameterizations, and 1-species networks in Section 3. In Section 4, we consider computational aspects of assessing ACR, introduce an ideal-decomposition algorithm for establishing ACR, and present several examples which illustrate the limitations to sufficiency and necessity of the presented conditions in establishing ACR. In Section 5, we present numerical methods to detect or to preclude ACR. Section 6 investigates methods for detecting “zero-divisor ACR”, and we end with a discussion in Section 7.
2. Mass-action systems and ACR
This section introduces chemical reaction networks (Section 2.1), mass-action systems (Section 2.2), their steady states (Section 2.3), and ACR (Section 2.4). We also recall several facts about ideals in polynomial rings, which we use later (Section 2.5).
2.1. Chemical reaction networks
Definition 2.1.
[…]
  1. (1)
[…]
  2. (2)
    a set $\mathcal{C}=\{y_{1},y_{2},\dots,y_{p}\}$ of complexes (finite nonnegative-integer combinations of the species), and
  3. (3)
    a set of reactions, ordered pairs of the complexes: $\mathcal{R}\subseteq(\mathcal{C}\times\mathcal{C})\smallsetminus\{(y,y)\mid y\in\mathcal{C}\}$.
A network can be viewed as a directed graph whose nodes are complexes and whose edges correspond to the reactions. Next, writing the $i$-th complex as $y_{i1}A_{1}+y_{i2}A_{2}+\cdots+y_{in}A_{n}$ (where $y_{ij}\in\mathbb{Z}_{\geq 0}$ for $j=1,2,\dots,n$), we introduce the following monomial:
$x^{y_{i}}\,\,\,:=\,\,\,x_{1}^{y_{i1}}x_{2}^{y_{i2}}\cdots x_{n}^{y_{in}}~{}.$
(By convention, the zero complex yields the monomial $x^{(0,\dots,0)}=1$.) The vectors $y_{i}$ define the rows of a $p\times n$-matrix of nonnegative integers, which we denote by $Y=(y_{ij})$. Next, the unknowns $x_{1},x_{2},\ldots,x_{n}$ represent the concentrations of the $n$ species in the network, and we regard them as functions $x_{i}(t)$ of time $t$.
For a reaction $y_{i}\to y_{j}$ from the $i$-th complex to the $j$-th complex, the reaction vector $y_{j}-y_{i}$ encodes the net change in each species that results when the reaction takes place. The stoichiometric subspace is the vector subspace of $\mathbb{R}^{n}$ spanned by the reaction vectors $y_{j}-y_{i}$, and we will denote this space by ${S}$:
(1) ${S}~{}:=~{}\mathbb{R}\{y_{j}-y_{i}\mid y_{i}\to y_{j}~{}{\rm is~{}in~{}}\mathcal{R}\}.$
Networks appearing in our examples have only a few species, and so we denote the species by $A,B,C,\dots$ instead of $X_{1},X_{2},X_{3}\dots$.
Example 2.2.
Consider the network $G=\{A+B\to 3A+C\}$, which consists of a single reaction (with 3 species and 2 complexes). We have $y_{2}-y_{1}=(2,-1,1)$, which means that with each occurrence of the reaction, two units of $A$ and one of $C$ are produced, while one unit of $B$ is consumed. This vector $(2,-1,1)$ spans the stoichiometric subspace ${S}$.
2.2. Mass-action systems
The dynamical systems considered in this work come from assigning mass-action kinetics to a chemical reaction network. According to the law of mass-action, the rate of each chemical reaction is directly proportional to the product of the concentrations of the reactants. Therefore, we associate to each reaction $y_{i}\to y_{j}$ of a network, a positive parameter $\kappa_{ij}$, the rate constant of the reaction.
Definition 2.3.
[…]
(2) $\frac{dx}{dt}\quad=\quad\sum_{y_{i}\to y_{j}~{}{\rm is~{}in~{}}\mathcal{R}}\kappa_{ij}x^{y_{i}}(y_{j}-y_{i})\quad=:\quad f_{\kappa}(x)~{}=~{}(f_{\kappa}(x)_{1},f_{\kappa}(x)_{2},\dots,f_{\kappa}(x)_{n})~{}.$
In this article, we treat the rate constants $\kappa_{ij}$ as (positive) unknowns in order to analyze the entire family of dynamical systems that arise from a given network, as the $\kappa_{ij}$’s vary.
By construction, the vector $\frac{dx}{dt}$ in (2) lies in ${S}$ for all time $t$. Therefore, a trajectory $x(t)$ beginning at a positive vector $x(0)=x^{0}\in\mathbb{R}^{n}_{>0}$ remains in the stoichiometric compatibility class, which we denote by
(3) $\mathcal{P}~{}:=~{}(x^{0}+{S})\cap\mathbb{R}^{n}_{\geq 0}~{},$
for all positive time. In other words, $\mathcal{P}$ is forward-invariant with respect to the dynamics (2).
Example 2.4.
The following is a “generalized Shinar–Feinberg network” [52]:
(4) $\{B\to A,~{}2A+B\to A+2B\}~{}.$
The mass-action ODEs (2) arising from this network are as follows:
(5) $\frac{dx_{A}}{dt}~{}$ $=~{}\kappa_{1}x_{B}-\kappa_{2}x_{A}^{2}x_{B}$
$\frac{dx_{B}}{dt}~{}$ $=~{}-\kappa_{1}x_{B}+\kappa_{2}x_{A}^{2}x_{B}~{},$
where $\kappa_{1}$ and $\kappa_{2}$, respectively, are the rate constants for the first and second reactions in (4). The stoichiometric subspace is spanned by the vector $(1,-1)$, so the stoichiometric compatibility classes are the line segments defined by $c\in\mathbb{R}_{>0}$, as follows:
(6) $\mathcal{P}_{c}~{}=~{}\{(x_{A},~{}x_{B})\in\mathbb{R}^{2}_{\geq 0}\mid x_{A}+x_{B}=c\}~{}.$
2.3. Steady states and related ideals
Here we define steady states and the polynomial ideals that they generate.
Definition 2.5.
Consider a mass-action system $(G,\kappa)$ with $n$ species.
  1. (1)
    A steady state is a nonnegative concentration vector $x^{*}\in\mathbb{R}_{\geq 0}^{n}$ at which the ODEs (2) vanish: $f_{\kappa}(x^{*})=0$. We distinguish between positive steady states $x^{*}\in\mathbb{R}^{n}_{>0}$ and boundary steady states $x^{*}\in\left(\mathbb{R}^{n}_{\geq 0}\smallsetminus\mathbb{R}^{n}_{>0}\right)$.
  2. (2)
    The positive steady-state locus is the set of all positive steady states.
  3. (3)
    The steady-state ideal is the ideal in the polynomial ring $\mathbb{R}[x_{1},x_{2},\ldots,x_{n}]$ generated by the right-hand sides of the ODEs in (2):
    $I(G,\kappa)~{}:=~{}\langle f_{\kappa}(x)_{1},~{}f_{\kappa}(x)_{2},~{}\dots,~{}f_{\kappa}(x)_{n}\rangle~{}.$
  4. (4)
    The ideal of the positive steady-state locus is defined as follows:
    $>0]{I(G,\kappa)}~{}:=~{}\left\{h\in\mathbb{R}[x_{1},x_{2},\ldots,x_{n}]\ |\ h(x)=0\text{ for all positive steady states }x\right\}~{}.$
We recall the definition of the radical of an ideal $I$ of a ring $R$:
(7) $\sqrt{I}~{}:=~{}\left\{h\in R\ |\ h^{m}\in I\text{ for some }m\in{\mathbb{Z}}_{>0}\right\}~{}.$
Remark 2.6.
By construction, we have $I(G,\kappa)\subseteq\sqrt[>0]{I(G,\kappa)}$. This containment of ideals can be refined, yielding three “levels” of “steady-state invariants” [21, equation (7)]:
(8) $I(G,\kappa)~{}\subseteq~{}\sqrt{I(G,\kappa)}~{}\subseteq~{}\sqrt[>0]{I(G,\kappa)}~{}.$
Here, $\sqrt{I(G,\kappa)}$ denotes the radical of the steady-state ideal. The fact that the containments in (8) are, in general, strict, was noted by Dickenstein [21] [^1] We make a small clarification: Dickenstein considered polynomials vanishing on all steady states, whereas we consider only positive steady states. As a result, the ideal Dickenstein called the “positive real radical of the steady-state ideal” is closely related to, but not the same as, our ideal $>0]{I(G,\kappa)}$. and is the reason behind many of the “warning” examples shown later in Section 4.4.
Example 2.7.
Consider the network $G=\{2A\to A,~{}3A\to 5A,~{}4A\to 3A\}$. When all three rate constants are taken to be 1 (that is, $\kappa^{*}=(1,1,1)$), the resulting mass-action ODE is the following:
$\frac{dx_{A}}{dt}~{}=~{}-x_{A}^{4}+2x_{A}^{3}-x_{A}^{2}~{}=~{}-x_{A}^{2}(x_{A}-1)^{2}~{}.$
(Notably, $x_{A}^{*}=1$ is the unique positive steady state.) The ideals in (8) are thus as follows:
$I(G,\kappa^{*})~{}=~{}\langle x_{A}^{2}(x_{A}-1)^{2}\rangle~{}\subsetneq~{}\sqrt{I(G,\kappa^{*})}~{}=~{}\langle x_{A}(x_{A}-1)\rangle~{}\subsetneq~{}\sqrt[>0]{I(G,\kappa^{*})}~{}=~{}\langle x_{A}-1\rangle~{}.$
Next, we move from systems $(G,\kappa)$ to networks $G$ by viewing $\kappa$ as unknown.
Definition 2.8.
Let $G$ be a chemical reaction network with $n$ species and $r$ reactions.
  1. (1)
    $G$ is multistationary if there exists positive rate constants $\kappa\in\mathbb{R}^{r}_{>0}$ such that the resulting mass-action kinetics system (2) admits two or more positive steady states in some stoichiometric compatibility class (3).
  2. (2)
    The positive steady-state/rate locus is the set of all pairs of positive steady states and corresponding rate constants:
    :=$\mathcal{L}~{}:=~{}\left\{(x,\kappa)\in\mathbb{R}^{n}_{>0}\times\mathbb{R}^{r}_{>0}\mid x\text{ is a steady state of the mass-action system }(G,\kappa)\right\}~{}.$
  3. (3)
    The ideal of the positive steady-state/rate locus is:
    $I(\mathcal{L})~{}:=~{}\left\{h\in\mathbb{R}[x_{1},x_{2},\ldots,x_{n};\kappa_{ij}]\mid h(x,\kappa)=0\text{ for all }(x,\kappa)\in\mathcal{L}\right\}~{}.$
Example 2.9 (Example 2.4, continued).
For the network $G=\{B\overset{\kappa_{1}}{\to}A,~{}2A+B\overset{\kappa_{2}}{\to}A+2B\}$, it is straightforward to compute the positive steady-state/rate locus:
(9) =$\mathcal{L}~{}=~{}\left\{(x,\kappa)\in\mathbb{R}^{2}_{>0}\times\mathbb{R}^{2}_{>0}\mid x_{A}=\sqrt{\kappa_{1}/\kappa_{2}}\right\}~{}.$
It follows that the ideal of the positive steady-state/rate locus is the following:
(10) $I(\mathcal{L})~{}=~{}\langle\kappa_{1}-\kappa_{2}x_{A}^{2}\rangle~{}.$
We observe that, for every choice of positive rate constants $\kappa^{*}=(\kappa^{*}_{1},\kappa^{*}_{2})\in\mathbb{R}^{2}_{>0}$, every positive steady state $(x_{A}^{*},x_{B}^{*})$ of the mass-action system $(G,\kappa^{*})$ has the same value of $x_{A}^{*}$, namely, $x^{*}_{A}=\sqrt{\kappa^{*}_{1}/\kappa^{*}_{2}}$. In the following subsection, we give a name to this property.
Remark 2.10.
Any positive part of a variety is equal to the positive steady-state locus of some mass-action system, because we can just multiply all the equations by $x_{1}x_{2}\dots x_{s}$ so that the resulting equations are mass-action [21, §2]. Indeed, many of the examples in this work were generated through such a trick.
2.4. Absolute concentration robustness
In this subsection, we recall the definition of ACR for mass-action systems and networks (Definitions 2.11 and 2.16). We also relate ACR to steady-state ideals (Propositions 2.15 and 2.19).
2.4.1. ACR for mass-action systems
The following definition is due to Shinar and Feinberg [68].
Definition 2.11 (ACR for mass-action systems).
A mass-action system $(G,\kappa)$ has absolute concentration robustness (ACR) in species $X_{i}$ if, for every positive steady state $x$, the value of $x_{i}$ is the same. This value of $x_{i}$ is the ACR-value.
For simplicity, we say that $(G,\kappa)$ has ACR if it has ACR is some species.
Definition 2.12 (Vacuous ACR).
A mass-action system $(G,\kappa)$ has vacuous ACR if it has no positive steady states (or, equivalently, the ideal of the positive steady-state locus is the improper ideal $\langle 1\rangle$).
An example of a system with vacuous ACR arises from the network $G=\{A\to B\}$.
Example 2.13 (Example 2.7 continued).
We saw earlier that the mass-action system given by the network $G=\{2A\to A,~{}3A\to 5A,~{}4A\to 3A\}$ and rate constants $\kappa^{*}=(1,1,1)$ has a unique positive steady state (namely, $x_{A}^{*}=1$). It follows that this system has ACR (in $A$).
Example 2.14 (Example 2.9 continued).
We observed above that, for $G=\{B\to A,~{}2A+B\to A+2B\}$, the mass-action system $(G,\kappa^{*})$ arising from any choice of positive rate constants $\kappa$ has ACR in $A$.
The following result, which was observed in [42], relates ACR with the ideal of the positive steady state locus (from Definition 2.5).
Proposition 2.15 (ACR and ideals).
A mass-action system $(G,\kappa)$ has ACR in species $X_{i}$ if and only if $x_{i}-\alpha$, for some $\alpha>0$, is in the ideal of the positive steady state locus of $(G,\kappa)$.
Proof.
If $(G,\kappa)$ has no positive steady states (that is, $>0]{I(G,\kappa)}=\langle 1\rangle$), then $x_{i}-1$ is in $>0]{I(G,\kappa)}$ and $(G,\kappa)$ vacuously has ACR (recall Definition 2.12). Now assume that $(G,\kappa)$ has at least one positive steady state. Then $(G,\kappa)$ has ACR in $X_{i}$ with ACR-value $\alpha>0$ if and only if $x^{*}_{i}=\alpha$ for every positive steady state $x^{*}$. The result now follows directly. ∎
2.4.2. ACR for networks
If $(G,\kappa)$ has ACR, then does ACR persist when $\kappa$ is replaced by another choice of rate constants? This question motivates the following definition.
Definition 2.16 (ACR for networks).
A network $G$ with $r$ reactions has:
  1. (i)
    the capacity for ACR if there exist a species $X_{i}$ and some $\kappa\in\mathbb{R}_{>0}^{r}$ such that $(G,\kappa)$ has ACR in $X_{i}$;
  2. (ii)
    unconditional ACR (or, simply, ACR) if there exists a species $X_{i}$ such that for all $\kappa\in\mathbb{R}_{>0}^{r}$, the mass-action system $(G,\kappa)$ has ACR in $X_{i}$.
Notice that networks with (unconditional) ACR automatically have the capacity for ACR. These ideas are related to “hybrid robustness” [19, pg. 889].
Example 2.17 (Example 2.14 continued).
From our earlier observations, we see that the network $\{B\to A,~{}2A+B\to A+2B\}$ has (unconditional) ACR (in species $A$, but not $B$).
Example 2.18 (Example 2.13 continued).
We saw that the network $G=\{2A\to A,~{}3A\to 5A,~{}4A\to 3A\}$ has the capacity for ACR. We now claim that $G$ does not have (unconditional) ACR. To see this, consider the rate constants $\kappa^{*}=(2,1.5,1)$. The resulting ODE is
$\frac{dx_{A}}{dt}~{}=~{}-x_{A}^{4}+3x_{A}^{3}-2x_{A}^{2}~{}=~{}-x_{A}^{2}(x_{A}-1)(x_{A}-2)~{},$
and so there are two positive steady states, $x_{A}^{*}=1$ and $x_{A}^{*}=2$. Hence, $(G,\kappa^{*})$ does not have ACR, which verifies the claim.
Recall that Proposition 2.15 pertained to ACR in mass-action systems (i.e., with specialized rate constants). We next state a version of that result for networks (i.e., with symbolic rate constants). Note that while Proposition 2.15 gives an equivalence between ACR and a condition on a certain ideal, the following result asserts only one such implication.
Proposition 2.19 (ACR and $I(\mathcal{L})$).
Let $X_{i}$ be a species of a reaction network $G$. Let $I(\mathcal{L})$ denote the ideal of the positive steady-state/rate locus of $G$. If $x_{i}-\alpha\in I(\mathcal{L})$, for some $\alpha\in\mathbb{R}[\kappa_{ij}\mid i\to j\mathrm{~{}is~{}a~{}reaction~{}of~{}}G]$, then $G$ has ACR in $X_{i}$.
Proof.
Assume there exists $\alpha\in\mathbb{R}[\kappa_{ij}\mid i\to j\mathrm{~{}is~{}a~{}reaction~{}of~{}}G]$ such that $x_{i}-\alpha\in I(\mathcal{L})$. Fix $\kappa^{*}\in\mathbb{R}^{r}_{>0}$, where $r$ is the number of reactions.
If the mass-action system $(G,\kappa^{*})$ has no positive steady states, then the system vacuously has ACR. Now assume that $(G,\kappa^{*})$ has a positive steady state $x^{*}$. Then, $x_{i}-\alpha\in I(\mathcal{L})$ implies that $x^{*}_{i}=\alpha|_{\kappa=\kappa^{*}}$. Thus, $(G,\kappa^{*})$ has ACR in $X_{i}$ with ACR-value $\alpha|_{\kappa=\kappa^{*}}$. ∎
The converse of Proposition 2.19 is false, as the following example shows.
Example 2.20 (Example 2.17 continued).
We noted earlier that the network $\{B\to A,~{}2A+B\to A+2B\}$ has ACR in $B$, but the ideal $I(\mathcal{L})$, in (10), does not contain a polynomial of the form $x_{i}-\alpha$ with $\alpha\in\mathbb{R}[\kappa_{1},\kappa_{2}]$.
The final example in this section previews the relationship between multistationarity and ACR, as well as the strong dependence of these properties on the choice of rate constants.
Example 2.21.
Consider the following one-species network:
[…]
where for convenience we are restricting our attention to a two-dimensional space of rate constants, as shown in the network above. The resulting ODE is
$\frac{dx_{A}}{dt}$ $~{}=~{}x_{A}^{3}-(\alpha+\beta)x_{A}^{2}+3x_{A}^{2}-(2\alpha+\beta)x_{A}+(3+\alpha\beta)x_{A}$
$~{}=~{}x_{A}(x_{A}-(\alpha-1))(x_{A}-(\beta-2))~{}.$
Thus, the ACR properties for this network depend on the rate constants, as follows:
  • If $\alpha>1$, $\beta>2$, and $\alpha\neq\beta-1$, then there are two positive steady states (namely, $x^{*}_{A}=\alpha-1$ and $x^{**}_{A}=\beta-2$) and so there is no ACR.
  • If $\alpha>1$, $\beta>2$, and $\alpha=\beta-1$, then there is a unique positive steady state ($x_{A}^{*}=\alpha-1$) and so there is ACR.
  • If $\alpha\leq 1$ and $\beta\leq 2$, then there are no positive steady states, and so there is vacuous ACR.
  • If $\alpha>1$ and $\beta\leq 2$, then there is a unique positive steady state ($x^{*}_{A}=\alpha-1$), and hence there is ACR.
  • If $\alpha\leq 1$ and $\beta>2$, then there is a unique positive steady state ($x^{*}_{A}=\beta-2$), and so there is ACR.
Visually, this is shown here:
[…]
We see that ACR in this system can be quite unstable. For instance, near $(\alpha,\beta)=(1,2)$, some parameter values yield ACR, while others yield no ACR or vacuous ACR. Also, in the ACR region with ACR-value $x_{A}=\beta-2$, when $\alpha$ is perturbed within the interval $(0,1)$ the ACR-value is perturbed; while in the ACR region with ACR-value $x_{A}=\alpha-1$, perturbing $\beta$ within the interval $(0,2)$ preserves the ACR-value.
2.5. Ideals in polynomial rings
We end this section by recalling several concepts related to general ideals (in polynomial rings): saturations, zero divisors, real varieties, and real radicals. We use these ideas in later sections (for steady-state ideals and other ideals).
Definition 2.22.
Let $h\in\mathbb{F}[x_{1},x_{2},\dots,x_{n}]$, where $\mathbb{F}$ is a characteristic-zero field, and let $J$ be an ideal of $\mathbb{K}[x_{1},x_{2},\dots,x_{n}]$. The saturation of $J$ with respect to $h$ is the following ideal:
$(J:h^{\infty})~{}:=~{}\{g\in\mathbb{F}[x_{1},x_{2},\dots,x_{n}]\mid h^{k}g\in J\;\text{for some}\;k\in\mathbb{Z}_{\geq 0}\}.$
Remark 2.23.
For the saturations we consider in this work, the polynomial $h$ is the monomial $\mathfrak{m}$ formed by the product of all the variables $x_{i}$ (that is, $\mathfrak{m}=x_{1}x_{2}\cdots x_{n}$). Roughly speaking, $(J:\mathfrak{m}^{\infty})$ allows us to “divide” the polynomials in $J$ by monomials in the variables. Hence, the zeros of $J$ with nonzero coordinates are the zeros of $(J:\mathfrak{m}^{\infty})$ with nonzero coordinates. Moreover, computing saturation is a standard task accomplished by Gröbner bases, which are implemented for instance in any computer algebra system.
Next, we turn to zero divisors.
Definition 2.24.
Let $I$ be an ideal in a ring $R$. If $fg\in I$, where $f,g\in R\smallsetminus I$, then $f$ and $g$ are zero divisors of $I$.
Remark 2.25.
In Definition 2.24, we use the term “zero divisor” because $f$ is a zero divisor of $I$ if and only if $f+I$ is a zero divisor of the quotient ring $R/I$.
Remark 2.26 (Checking zero divisors).
A polynomial $f$ is a zero divisor of an ideal $I$ in a polynomial ring, if and only if the saturation ideal $I:(f)$ is strictly larger than $I$. Checking this is easy using computer algebra software (as mentioned in Remark 2.23, Gröbner bases can be used to compute saturations – and the same is true for testing equality of ideals).
Finally, we consider real varieties and real radicals.
Definition 2.27.
Let $J$ be an ideal in the polynomial ring $\mathbb{Q}[x_{1},x_{2},\dots,x_{n}]$. The real radical of $J$ is the following ideal:
$\mathbb{R}]{J}~{}:=~{}\left\{g\in{\mathbb{Q}}[x_{1},x_{2},\dots,x_{n}]~{}|~{}\exists m,\ell\in\mathbb{Z}_{\geq 0},~{}\exists h_{1},\dots,h_{\ell}\in{\mathbb{Q}}[x_{1},x_{2},\dots,x_{n}],\;g^{2m}+\sum_{i=1}^{\ell}h_{i}^{2}\in J\right\}~{}.$
The real variety of $J$ is:
$V_{\mathbb{R}}(J)~{}:=~{}\left\{z\in\mathbb{R}^{n}\mid f(z)=0\text{ for all }f\in J\right\}.$
The real radical and real variety are related by the next result, the Real Nullstellensatz (specifically, the version in which the coefficient field is $\mathbb{Q}$). This result is well known; for instance, it follows from [50, Theorem 1.11] [^2] A proof of Proposition 2.28 is in Appendix A.. Neuhaus [60] credits the result to Dubois [23], Krivine [44], and Risler [65].
Proposition 2.28 (Real Nullstellensatz).
If $J$ is an ideal in the polynomial ring $\mathbb{Q}[x_{1},x_{2},\dots,x_{n}]$, then the real radical of $J$ equals the vanishing ideal of the real variety of $J$:
$\mathbb{R}]{J}~{}:=~{}\left\{f\in\mathbb{Q}[x_{1},x_{2},\dots,x_{n}]\mid f(z)=0\text{ for all }z\in V_{\mathbb{R}}(J)\right\}~{}.$
3. Multistationarity, Steady-state parameterizations, and 1-species networks
In this section, we analyze the relationship between ACR and multistationarity (Section 3.1). We also explain how to assess ACR for systems that admit a steady-state parameterization (Section 3.2) and those having only one species (Section 3.3).
3.1. Multistationarity
In this subsection, we analyze the relationship between ACR and multistationarity. For networks with only one species, ACR in one species is trivially the same as ACR in all species, and having ACR is equivalent to being non-multistationary [52, Proposition 4.2]. Here we show that one implication of this result holds in general (Lemma 3.2), while the full result generalizes for networks with no linear conservation laws or, equivalently, when the stoichiometric subspace is all of $\mathbb{R}^{n}$ (Proposition 3.3). Finally, we show that outside of the constraints imposed by these two results, all other possibilities for ACR and multistationarity are realized (see Table 1).
Remark 3.1.
The relationship between ACR and multistationarity was also explored in prior works, in the context of randomly generated reaction networks [39] and for the purpose of determining the minimum numbers of species, complexes, and reaction needed for ACR and multistationarity to coexist [41].
Lemma 3.2.
If a reaction network $G$ has ACR in every species, then $G$ is not multistationary.
Proof.
This is immediate from the definitions of ACR and multistationary. ∎
Proposition 3.3.
Let $G$ be a network with $n$ species. If the stoichiometric subspace of $G$ is $\mathbb{R}^{n}$, then $G$ has ACR in every species $X_{i}$ if and only if $G$ is not multistationary.
Proof.
The implication $\Rightarrow$ is Lemma 3.2. For $\Leftarrow$, assume that $G$ is non-multistationary. Fix a choice of positive rate constants $\kappa$. If the mass-action system $(G,\kappa)$ has no positive steady states, then it has vacuous ACR in every species (recall Definition 2.12). The remaining case is when $(G,\kappa)$ has a unique positive steady state $x^{*}$ (because $G$ is non-multistationary and $\mathcal{P}=\mathbb{R}^{n}_{>0}$ is the unique stoichiometric compatibility class (3)). In this case, $(G,\kappa)$ has ACR in every species $X_{i}$ (with corresponding ACR-value $x_{i}^{*}$). ∎
Remark 3.4.
Proposition 3.3 pertains to ACR for networks with full-dimensional stoichiometric subspace. This result therefore complements recent results of Meshkat, Shiu, and Torres characterizing ACR in networks with one-dimensional stoichiometric subspace and those with up to two reactions [52].
Example 3.5.
Consider the network $G=\{0\to A\to B\to 0\}$. It is easy to check that the stoichiometric subspace is $\mathbb{R}^{2}$. Also, for all choices of positive rate constants, there is a unique positive steady state (one can check this by hand or by applying the Deficiency Zero Theorem [26]). Therefore, $G$ is not multistationary and so, by Proposition 3.3, has ACR in every species.
Lemma 3.2 and Proposition 3.3 involve the property of having ACR in all species. In general, we are more interested in ACR in some species. We now show that we can realize all possibilities involving three properties – ACR in all species, ACR in some species, and multistationarity – outside of the constraints arising from Lemma 3.2, Proposition 3.3, and the fact that having ACR in all species implies ACR in some species. We summarize this information in Table 1, which is inspired by [37, Table 1].
$S=\mathbb{R}^{n}$? Multistationary? ACR in all? ACR is some? Example
yes yes no yes Example 3.6
yes yes no no [52, Example 2.32]
yes no yes yes Example 3.5
no yes no yes [52, Example 2.12]
no yes no no Example 3.7
no no no yes Example 2.17
no no no no Example 3.11
no no yes yes Example 3.8
Table 1. Possible properties of reaction networks – having full-dimensional stoichiometric subspace ($S=\mathbb{R}^{n}$), being multistationary, having ACR in all species, and having ACR in some species – and corresponding examples showing such networks. Among the example networks in this table with ACR, only Example 3.8 has vacuous ACR (no positive steady states); we conjecture that no example with non-vacuous ACR exists for this line of the table.
Example 3.6.
Consider the following network:
$G~{}=~{}\left\{A+B\overset{\kappa_{1}}{\underset{\kappa_{2}}{\leftrightarrows}}B\overset{\kappa_{3}}{\underset{\kappa_{4}}{\leftrightarrows}}0~{},~{}2B\overset{\kappa_{5}}{\underset{\kappa_{6}}{\leftrightarrows}}3B\right\}~{}.$
It is straightforward to check that $G$ has no conservation laws and that for any positive steady state $(x_{A}^{*},x_{B}^{*})$, we have $x_{A}^{*}=\kappa_{2}/\kappa_{1}$. Thus, $G$ has ACR in species $A$. Also, $x_{B}^{*}$ can take up to three positive-steady-state values. For instance, if $\kappa_{3}=11$, $\kappa_{4}=\kappa_{5}=6$, and $\kappa_{6}=1$, then $x_{B}^{*}$ takes three values (namely, $x_{B}^{*}=1$, 2, and 3), yielding three positive steady states $(x_{A}^{*}=\kappa_{2}/\kappa_{1},~{}x_{B}^{*})$. Hence, $G$ is not multistationary and does not have ACR in $B$.
Example 3.7.
The network $G=\left\{3A\overset{\kappa_{1}}{\underset{\kappa_{2}}{\rightleftarrows}}2A+B~{},~{}A+2B\overset{\kappa_{3}}{\underset{\kappa_{4}}{\rightleftarrows}}3B\right\}$ is known to be multistationary [34, 71]. The stoichiometric subspace is one-dimensional. It is straightforward to check that the stoichiometric compatibility class $\mathcal{P}=\{(x_{A},x_{B})\in\mathbb{R}^{2}_{\geq 0}\mid x_{A}+x_{B}=4\}$ contains three positive steady states, namely, $(1,3)$, $(2,2)$, and $(3,1)$. The network $G$ therefore does not have ACR.
Example 3.8.
The network $\{A\to B\}$ admits no positive steady states. Hence, it is non-multistationary and has vacuous ACR in both species. Also, the stoichiometric subspace is one-dimensional (in $\mathbb{R}^{2}$).
3.2. Steady-state parameterizations
For many networks arising in biology, the positive steady-state locus can be parameterized [15, 27, 36, 64, 74]. More precisely, we say that a mass-action system $(G,\kappa)$ with $n$ species admits a positive steady-state parameterization if there exists a function:
(11) $\phi:~{}\mathbb{R}^{T}_{>0}~{}$ >$\to~{}\mathbb{R}^{n}_{>0}$
(for some $T>0$) such that the image equals the set of all positive steady states of $(G,\kappa)$. When there is a steady-state parameterization, assessing ACR is easy, as follows.
Proposition 3.9.
Let $(G,\kappa)$ be a mass-action system that admits a positive steady-state parameterization, as in (11). Then $(G,\kappa)$ has ACR in species $X_{i}$ if and only if $\phi_{i}$ is constant.
Proof.
This follows from the definitions of ACR and positive steady-state parameterization. ∎
Example 3.10 (Example 2.20 continued).
We revisit the network $\{B\to A,~{}2A+B\to A+2B\}$. From (9), we see that for any choice of positive rate constants $\kappa_{1},\kappa_{2}$, the resulting mass-action system has the following positive steady-state parameterization:
$\phi:~{}\mathbb{R}_{>0}~{}$ >$\to~{}\mathbb{R}^{2}_{>0}$
$b~{}$ $\mapsto~{}(\sqrt{\kappa_{2}/\kappa_{1}}~{},~{}b)~{}.$
As $\phi_{1}$ is constant, Proposition 3.9 confirms, as observed earlier, ACR in species $A$.
Example 3.11.
Consider the following network, the reversible Michaelis-Menten mechanism:
$G~{}=~{}\left\{S_{0}+K\overset{\kappa_{1}}{\underset{\kappa_{2}}{\rightleftarrows}}S_{0}K\overset{\kappa_{3}}{\underset{\kappa_{4}}{\rightleftarrows}}S_{1}+K\right\}~{}.$
Let $x_{1},x_{2},x_{3},x_{4}$ denote the species concentrations for $S_{0},K,S_{0}K,S_{1}$, respectively. It is straightforward to check that (for any choice of $\kappa\in\mathbb{R}^{4}_{>0}$) the following is a positive steady-state parameterization (see [15] for details):
$\phi:~{}\mathbb{R}^{2}_{>0}~{}$ >$\to~{}\mathbb{R}^{4}_{>0}$
$(x_{2},x_{4})~{}$ $\mapsto~{}\left(\frac{\kappa_{2}\kappa_{4}}{\kappa_{1}\kappa_{3}}x_{4},~{}x_{2},~{}\frac{\kappa_{4}}{\kappa_{3}}x_{2}x_{4},~{}x_{4}\right)~{}.$
None of the coordinates $\phi_{i}$ is constant (each depends on at least one $x_{j}$). Hence, Proposition 3.9 implies that $G$ does not have ACR in any species.
Another property of this network is that, for all positive rate constants, the mass-action system yields a unique positive steady state in each (two-dimensional) stoichiometric compatibility class. This can be verified directly or by applying the Deficiency Zero Theorem [26]. It follows that $G$ is not multistationary.
3.3. One-species networks and systems
For networks having only one species, ACR has already been classified: Such a network has ACR if and only if it is non-multistationary, which in turn is equivalent to a combinatorial criterion on its “arrow diagram” [52, Proposition 4.2].
Next, we assume that the rate constants are fixed, that is, we consider one-species mass-action systems. In this setting, ACR means that the (univariate polynomial) right-hand side of the (unique) ODE has at most one positive root. This property can be checked using the classical theory of Sturm sequences, and we summarize this fact as follows.
Lemma 3.12.
There is an algorithm with the following input and output:
  • Input: A nonzero, univariate polynomial $g$ with rational coefficients (that is, $0\neq g\in\mathbb{Q}[x_{1}]$).
  • Output: The number of distinct positive real roots of $g$.
Observe that Lemma 3.12 concerns polynomials whose coefficients are rational numbers. The reason we avoid irrational coefficients is to circumvent issues regarding whether we can compute over the real numbers. Accordingly, in applying Lemma 3.12, we restrict our attention to systems in which every rate constant is rational, as follows.
Proposition 3.13 (Deciding ACR for one-species systems).
There is an algorithm with the following input and output:
  • Input: A mass-action system $(G,\kappa)$ with only one species and for which every rate constant is a rational number (that is, $\kappa\in\mathbb{Q}^{r}$, where $r$ is the number of reactions).
  • Output: “Vacuous ACR”, “Non-vacuous ACR”, or “No ACR”, depending on whether $(G,\kappa)$ has (respectively) vacuous ACR, non-vacuous ACR, or no ACR.
Proof.
We claim that the following steps yield an algorithm with the desired properties:
  1. (1)
    If $G$ has no reactions, then output “Vacuous ACR”.
  2. (2)
    Else, perform the algorithm from Lemma 3.12 on the right-hand side of the mass-action ODE for $(G,\kappa)$. Denote the output of that algorithm by $N$.
    1. (a)
      If $N=0$, output “Vacuous ACR”.
    2. (b)
      If $N=1$, output “Non-vacuous ACR”.
    3. (c)
      If $N\geq 2$, output “No ACR”.
For this algorithm, the proof of correctness follows easily from Lemma 3.12 and the definition of ACR and vacuous ACR. ∎
In the next section, we further discuss this assumption of rational-number rate constants, as well as other computational considerations.
4. ACR and computational considerations
In this section, we consider the computational aspects of the problem of deciding whether a system $(G,\kappa)$ has non-vacuous ACR. We first explain why, in many of our results, we assume that the rate constants are rational (Section 4.1). Next, we show that, in some cases, we can use a new ideal, which we call the positive-restriction ideal (Section 4.2), and also decompositions of its radical ideal (Section 4.3) to assess ACR. Finally, Section 4.4 serves as a warning, by showing that various sufficient conditions for ACR are not necessary.
4.1. On the assumption of rational-number rate constants
We have seen that it is sometimes convenient to assume that the rate constants are rational numbers (as in Proposition 3.13). We make this assumption again for some results in this section.
The reason for this assumption, in the previous section, was to avoid issues concerning whether and how we can compute over the real numbers. Another reason comes from the fact that, in real-life applications, we are interested in ACR (and other properties) that persist when rate constants are perturbed. In such situations, approximating real-number rate constants by rational numbers would suffice.
Nevertheless, in principle, the assumption of rational rate constants may be problematic. There are two main difficulties. The first is the issue of approximating the (actual) real-number coefficients by rational ones. The second arises when the ACR-value is irrational, and then this value either is approximated (and so propagates numerical errors) or is described symbolically (but possibly in a way that is not informative). Such difficulties are shown in the next examples.
Example 4.1.
[…]
$\frac{dx_{A}}{dt}~{}=~{}x_{A}^{2}-\pi x_{A}+3.1415x_{A}~{}=~{}x_{A}(x_{A}-(\pi-3.1415))~{},$
If we approximate $\pi$ by the rational number $3.1416$, or any number bigger than $3.1415$, then the resulting system has ACR. However, if we instead choose a number less than or equal to $3.1415$, there are no positive steady states and so the system has vacuous ACR. Hence, the precision of the approximation of the rate constant matters.
The example below shows the problem of dealing numerically with irrational-number roots.
Example 4.2.
[…]
$\frac{dx_{A}}{dt}$ $=~{}2-x_{A}^{2}~{},$
$\frac{dx_{B}}{dt}$ $=~{}x_{B}^{2}-x_{A}x_{B}+1.41x_{B}~{}=~{}x_{B}(x_{B}-(x_{A}-1.41))~{}.$
The elimination ideal $I\cap{\mathbb{Q}}[x_{A}]$, where $I$ is the steady-state ideal, is generated by $2-x_{A}^{2}$. Hence, this system has ACR with ACR-value $\sqrt{2}$. However, numerical problems may arise when checking whether ACR is vacuous or non-vacuous. Specifically, when we try to find the possible steady-state values of $x_{B}$, we have the following issue: as in Example 4.1, if we approximate $x_{A}=\sqrt{2}$ by any rational number less than or equal to $1.41$, we do not obtain any positive steady states and so we would conclude that the system has vacuous ACR. However, the system has non-vacuous ACR in both species, because the only positive steady state is $(x_{A}^{*},x_{B}^{*})=(\sqrt{2},\sqrt{2}-1.41)$.
4.2. Assessing ACR using the positive-restriction ideal
As explained in the prior subsection, the results in this subsection pertain to systems with rational rate constants. Specifically, we show that, for such systems, ACR can be detected from the real radical of an ideal we call the positive-restriction ideal (Proposition 4.7). This ideal expands the steady-state ideal (and resides in a polynomial ring with additional variables $z_{i}$) so that the resulting real variety forms a “cover” of the original positive steady states (see Lemma 4.5). The purpose of this ideal, therefore, is to restrict our attention to the positive steady states (rather than including boundary ones), hence the name “positive-restriction ideal”.
Definition 4.3.
Consider a mass-action system $(G,\kappa)$ with $n$ species and $r$ reactions, where $\kappa\in\mathbb{Q}_{>0}^{r}$. The positive-restriction ideal of $(G,\kappa)$ is the ideal in ${\mathbb{Q}}[x_{1},x_{2},\dots,x_{n},z_{1},z_{2},\dots,z_{n}]$ generated by the right-hand sides of the ODEs in (2), $f_{\kappa}(x)_{1},~{}f_{\kappa}(x)_{2},~{}\dots,~{}f_{\kappa}(x)_{n}$ (viewed here as polynomials in ${\mathbb{Q}}[x_{1},x_{2},\dots,x_{n},z_{1},z_{2},\dots,z_{n}]$) and the polynomials of the form $x_{i}z_{i}^{2}-1$:
(12) $J(G,\kappa)~{}:=~{}\langle f_{\kappa}(x)_{1},~{}f_{\kappa}(x)_{2},~{}\dots,~{}f_{\kappa}(x)_{n},~{}x_{1}z_{1}^{2}-1,~{}x_{2}z_{2}^{2}-1,~{}\dots,~{}x_{n}z_{n}^{2}-1\rangle~{}.$
Remark 4.4.
The use of the polynomials $x_{i}z_{i}^{2}-1$ is a version of the ‘Rabinowitsch trick’.
Lemma 4.5.
For a reaction system $(G,\kappa)$ with $n$ species and $\kappa\in\mathbb{Q}_{>0}^{r}$ (where $r$ is the number of reactions), the following projection (to the first $n$ coordinates) is a $2^{n}$-to-$1$ surjection from the real variety of the positive-restriction ideal to the positive steady-state locus:
$V_{\mathbb{R}}(J(G,\kappa))$ $~{}\overset{\phi}{\to}~{}V_{>0}(I(G,\kappa))$
$(x;z)$ $~{}\mapsto~{}x~{}.$
Proof.
We first show that the image of $\phi$ is contained in $V_{>0}(I(G,\kappa))$. Assume that $(x;z)\in V_{\mathbb{R}}(J(G,\kappa))$. Then $x_{i}z_{i}^{2}-1=0$ for all $i$, so $z_{i}\neq 0$ and $x_{i}>0$ for all $i$. The vector $x$ is also a root of the polynomials in $I(G,\kappa)$, since the generators of $I(G,\kappa)$ also belong to $J(G,\kappa)$. We conclude that $x\in V_{>0}(I(G,\kappa))$.
Next, let $x\in V_{>0}(I(G,\kappa))$. For all $i$, let $z_{i}:=\pm\sqrt{1/x_{i}}$. It follows that $(x;z)\in V_{\mathbb{R}}(J(G,\kappa))$. It is straightfoward to see that no other choices of $z_{i}$ yield $(x;z)\in V_{\mathbb{R}}(J(G,\kappa))$. So, $\phi$ is surjective and is $2^{n}$-to-$1$. ∎
Remark 4.6 (Viewing the positive-restriction ideal as a steady-state ideal).
For a mass-action system $(G,\kappa)$ with $n$ species, the positive-restriction ideal $J(G,\kappa)$ can be viewed as the steady-state ideal of the mass-action system obtained by adding the following $2n$ reactions to $(G,\kappa)$:
[…]
Next, we use Lemma 4.5 and the Real Nullstellensatz (Proposition 2.28) to extend Proposition 2.15, in the case when the rate constants – and also the ACR-value – are rational numbers.
Proposition 4.7 (Rational-number ACR and ideals).
Consider a mass-action system $(G,\kappa)$ with $\kappa\in\mathbb{Q}_{>0}^{r}$ (where $r$ is the number of reactions), and let $\alpha\in\mathbb{Q}_{>0}$. The following are equivalent:
  1. (1)
    $(G,\kappa)$ has ACR in species $X_{i}$ with ACR-value $\alpha$,
  2. (2)
    $x_{i}-\alpha$ is in the ideal of the positive steady state locus of $(G,\kappa)$,
  3. (3)
    $x_{i}-\alpha$ is in $\mathbb{R}]{J(G,\kappa)}$ (the real radical of the positive-restriction ideal).
Proof.
The equivalence of (1) and (2) is Proposition 2.15. Now we prove that (1) and (3) are equivalent. Let $\alpha>0$. By Lemma 4.5, every positive steady state $x^{*}$ satisfies $x_{i}^{*}=\alpha$ if and only if every $(x^{*};z^{*})$ in the real variety of the positive-restriction ideal satisfies $x^{*}_{i}=\alpha$, which in turn (by Proposition 2.28) is equivalent to the condition $(x_{i}-\alpha)\in\sqrt[\mathbb{R}]{J(G,\kappa)}$. This final equivalence we just asserted uses the hypothesis that $\alpha\in\mathbb{Q}_{>0}$; indeed, if instead the ACR-value $\alpha$ were irrational, then its irreducible polynomial (over $\mathbb{Q}$), rather than the linear polynomial $x_{i}-\alpha$, would be in $\mathbb{R}]{J(G,\kappa)}$. ∎
Remark 4.8.
In Proposition 4.7, it is assumed that not only are the rate constants rational, but also the ACR-value itself. This is not always the case (for instance, we saw in Example 2.9 a network in which the ACR-value has the form $\sqrt{\kappa_{1}/\kappa_{2}}$). Nevertheless, this property of rational-number ACR-value is common in the literature, and can be explained in part by the theory of “robust ratios” which, when applicable, gives a precise rational-number expression of the ACR-value in terms of the rate constants [75].
Remark 4.9.
When the ACR-value is irrational, Proposition 4.7 does not apply. Nevertheless, we can see from the proof of the proposition that sometimes ACR can be detected. Indeed, as in the end of the proof, if a univariate (in $x_{i}$), irreducible polynomial $g$ appears in the real radical of the positive-restriction ideal, and $g$ has a unique positive root $\alpha$, then there is ACR with ACR-value $\alpha$.
Proposition 4.7 reveals an advantage of considering rational-number rate constants: Detecting ACR reduces to being able to compute a real radical ideal – as long as the ACR-value is also rational.
In fact, real radicals can be computed effectively to some extent [60, 72], although the complexity of such algorithms is high. Moreover, new methods for computing real radicals have been developed recently [3, 66], but it is important to note that when these algorithms encounter irrational roots $\alpha$ of polynomials, either a numerical approximation of $\alpha$ is returned, or, for symbolic algorithms, the irreducible polynomial of $\alpha$. Similar drawbacks were shown earlier in Examples 4.1 and 4.2.
4.3. ACR via ideal decomposition
This subsection analyzes ACR by decomposing the positive-restriction ideal $J(G,\kappa)$ and then, using the notion of non-singular zeros, by restricting to certain components of $J(G,\kappa)$ (see Theorem 4.13). We also discuss issues involved in turning Theorem 4.13 into an algorithm for detecting ACR (when the ACR-value is rational).
We first introduce notation we use in this subsection. Given an ideal $P$ of ${\mathbb{Q}}[x_{1},x_{2},\dots,x_{n}]$, we let $\widetilde{P}$ denote the ideal of $\mathbb{R}[x_{1},x_{2},\dots,x_{n}]$ generated by the elements of $P$:
(13) $\widetilde{P}~{}:=~{}P\mathbb{R}[x_{1},x_{2},\dots,x_{n}].$
We now state a lemma that points in the direction we are pursuing.
Lemma 4.10.
If $J$ is an ideal of the ring $R:=\mathbb{Q}[x_{1},x_{2},\dots,x_{n}]$ that can be decomposed as $J=Q_{1}\cap Q_{2}$, where $Q_{1}$ and $Q_{2}$ are ideals in $R$ such that $V_{\mathbb{R}}(Q_{2})=\varnothing$, then $V_{\mathbb{R}}(J)=V_{\mathbb{R}}(Q_{1})$.
Proof.
This result is immediate since $V_{\mathbb{R}}(J)=V_{\mathbb{R}}(Q_{1})\cup V_{\mathbb{R}}(Q_{2})$. ∎
Next, we recall a criterion that can help us isolate certain components of the positive-restriction ideal. To state that result (Proposition 4.12 below), we must first recall several definitions (see [4, 8]). An ideal $I$ in a ring $R$ is prime if $xy\in I$ implies that $x\in I$ or $y\in I$. The dimension of an ideal $I$ of $\mathbb{R}[x_{1},x_{2},\dots,x_{n}]$ is the Krull dimension of the quotient ring $\mathbb{R}[x_{1},x_{2},\dots,x_{n}]/I$. If $I$ is generated by $n-d$ algebraically independent polynomials, then the dimension of $I$ is $d$ (see [45, Corollary 3.7]).
We now adapt to our context the definition of a non-singular zero, and subsequently recall a result that appears in the book of Bochnak, Coste, and Roy [8, Proposition 3.3.16].
Definition 4.11.
Let $P=\langle g_{1},g_{2},\dots,g_{t}\rangle$ be a prime ideal of $\mathbb{R}[x_{1},x_{2},\dots,x_{n}]$ of dimension $d$. A point $\mathbf{x}\in V_{\mathbb{R}}(P)$ is a non-singular zero of $P$ if $\mathrm{rank}\left(\left[\frac{\partial g_{i}}{\partial x_{j}}(\mathbf{x})\right]_{i,j}\right)=n-d$.
Proposition 4.12.
Let $P$ be a prime ideal of $\mathbb{R}[x_{1},x_{2},\dots,x_{n}]$. If $P$ has a non-singular zero, then $P=\sqrt[\mathbb{R}]{P}$.
From this proposition, we deduce the following result.
Theorem 4.13.
Consider a mass-action system $(G,\kappa)$ with $\kappa\in{\mathbb{Q}}_{0}^{r}$, and let $J(G,\kappa)$ be the positive-restriction ideal (12). Let $\alpha\in{\mathbb{Q}}_{>0}$. Assume the following:
  1. (1)
    $J(G,\kappa)=Q_{1}\cap Q_{2}$ for some ideals $Q_{1},Q_{2}$ of ${\mathbb{Q}}[x_{1},x_{2},\dots,x_{n},z_{1},z_{2},\dots,z_{n}]$ such that $V_{\mathbb{R}}(Q_{2})=\varnothing$, and
  2. (2)
    $Q_{1}=\overset{\ell}{\underset{i=1}{\bigcap}}P_{i}$ for some ideals $P_{i}$ of ${\mathbb{Q}}[x_{1},x_{2},\dots,x_{n},z_{1},z_{2},\dots,z_{n}]$ such that $V_{\mathbb{R}}(P_{i})\neq\varnothing$ (for all $i=1,2,\dots,\ell$).
Let $\widetilde{P}_{i}$, for all $i=1,2,\dots,\ell$, denote the corresponding ideals in $\mathbb{R}[x_{1},x_{2},\dots,x_{n},z_{1},z_{2},\dots,z_{n}]$, as in (13). Assume two additional main hypotheses:
  1. (3)
    $\widetilde{P}_{i}$ is a prime ideal of $\mathbb{R}[x_{1},x_{2},\dots,x_{n},z_{1},z_{2},\dots,z_{n}]$ (for all $i=1,\dots\ell$), and
  2. (4)
    $\widetilde{P}_{i}$ has a non-singular zero (for all $i=1,2,\dots,\ell$).
Then $(G,\kappa)$ has ACR in some species $X_{j}$ with ACR-value $\alpha$ if and only if $x_{j}-\alpha\in Q_{1}$.
Proof.
Let $J:=J(G,\kappa)$. By Proposition 4.7, it suffices to prove the equality $Q_{1}=\sqrt[\mathbb{R}]{J}$.
Hypothesis (1) and Lemma 4.10 together imply that $V_{\mathbb{R}}(J)=V_{\mathbb{R}}(Q_{1})$. Hence, by the Real Nullstellensatz (Proposition 2.28), we obtain the first equality here:
[…]
and the second equality follows from hypothesis (2) and the fact that the real radical of an intersection is the intersection of the real radicals [60, Lemma 2.2]. Next, equation (14) and the containment $Q_{1}\subseteq\sqrt[\mathbb{R}]{Q_{1}}$, together imply that $Q_{1}\subseteq\sqrt[\mathbb{R}]{J}$. Hence, using equation (14), it remains only to show the containment $\left(\overset{\ell}{\underset{i=1}{\bigcap}}\sqrt[\mathbb{R}]{P_{i}}\right)\subseteq Q_{1}$.
To verify this containment, we first note that $\mathbb{R}]{P_{i}}\subseteq\sqrt[\mathbb{R}]{\widetilde{P}_{i}}\cap{\mathbb{Q}}[x_{1},x_{2},\dots,x_{n},z_{1},z_{2},\dots,z_{n}]$ (for all $i=1,2,\dots,\ell$). Next, by hypotheses (34) and Proposition 4.12, we have $\mathbb{R}]{\widetilde{P}_{i}}=\widetilde{P}_{i}$. Finally, $\widetilde{P_{i}}\cap{\mathbb{Q}}[x_{1},x_{2},\dots,x_{n},z_{1},z_{2},\dots,z_{n}]=P_{i}$ [51, Theorem 7.5]. We conclude that $\left(\overset{\ell}{\underset{i=1}{\bigcap}}\sqrt[\mathbb{R}]{P_{i}}\right)\subseteq\overset{\ell}{\underset{i=1}{\bigcap}}P_{i}=Q_{1}$ (the equality is by hypothesis 2), so the desired containment holds. ∎
The hypotheses of Theorem 4.13 are, unfortunately, not easily checked. Indeed, in general, we cannot detect whether an ideal is prime over $\mathbb{R}$ or whether it has a non-singular zero. Nevertheless, when such computations are possible, the theorem tells us that if certain ideals $\widetilde{P}_{i}$ arising from a decomposition of $Q_{1}$ are prime over $\mathbb{R}$ and admit non-singular zeros, ACR with rational ACR-value is characterized by the presence of a linear polynomial $x_{i}-\alpha$ in $Q_{1}$. (See Steps 1 and 2 in the procedure below.)
But perhaps the most interesting case arises precisely when some of the ideals $\widetilde{P}_{i}$ lack non-singular zeros. In this case, all the real zeros of $J$ vanish at the polynomials that define the singular zeros of the corresponding component. Accordingly, we can add such polynomials to $J$ – without affecting its real variety – and instead analyze this larger ideal. This idea underlies the following procedure (in particular, Steps 3 and 4).
Procedure 4.14.
Detecting ACR by ideal decomposition
  • Input.
    An ideal $Q_{1}$ of ${\mathbb{Q}}[x_{1},x_{2},\dots,x_{n},z_{1},z_{2},\dots,z_{n}]$, such that $V_{\mathbb{R}}(Q_{1})=V_{\mathbb{R}}(J)$, where $J:=J(G,\kappa)$ is the positive-restriction ideal of a mass-action system $(G,\kappa)$.
  • Output.
    “ACR” or “Inconclusive”.
  • Step 1.
    Decompose $Q_{1}$ as $Q_{1}=\overset{\ell}{\underset{i=1}{\bigcap}}P_{i}$, where (for all $i=1,2,\dots,\ell$) $P_{i}$ is an ideal of the ring ${\mathbb{Q}}[x_{1},x_{2},\dots,x_{n},z_{1},z_{2},\dots,z_{n}]$ and $\widetilde{P}_{i}$, as in (13), is a prime ideal of the ring $\mathbb{R}[x_{1},x_{2},\dots,x_{n},z_{1},z_{2},\dots,z_{n}]$.
  • Step 2.
    If $\widetilde{P}_{i}$ has a non-singular zero for all $i=1,2,\dots,\ell$, then check whether $Q_{1}$ contains a polynomial of the form $x_{i}-\alpha$, where $\alpha\in\mathbb{Q}_{>0}$. If so, then output, “ACR in $X_{i}$ with ACR-value $\alpha$”.
  • Step 3.
    Otherwise, for every $P_{i}=\langle g_{1}^{i},g_{2}^{i},\dots,g_{t_{i}}^{i}\rangle$ of dimension $d_{i}$ with no non-singular zeros, add all the size-$(2n-d_{i})$ minors of the Jacobian matrix of $(g_{1}^{i},g_{2}^{i},\dots,g_{t_{i}}^{i})$ to $Q_{1}$. Call this new ideal $Q_{1}^{(1)}$.
  • Step 4.
    Repeat, if possible, Steps 1–3 for the ideal $Q_{1}=Q_{1}^{(1)}$.
By repeating Steps 1–4, we can in theory generate an ascending chain of ideals $Q_{1}\subseteq Q_{1}^{(1)}\subseteq Q_{1}^{(2)}\subseteq\dots\subseteq Q_{1}^{(j)}\subseteq\cdots$ that stabilizes, since the ring ${\mathbb{Q}}[x_{1},x_{2},\dots,x_{n},z_{1},z_{2},\dots,z_{n}]$ is Noetherian. The advantage of working with the ideal $Q_{1}^{(j)}$ instead of $J$ is that its dimension is smaller (and yet it has the same real variety).
For a better understanding of this theoretical procedure, and how to apply it to compute and detect ACR (when all steps can be done effectively), consider the following example:
Example 4.15.
[…]
$\frac{dx_{A}}{dt}$ $=~{}x_{A}^{3}+x_{A}x_{B}^{2}-2x_{A}^{2}-4x_{A}x_{B}+5x_{A}~{}=:f_{A}$
$=~{}x_{A}[(x_{A}-1)^{2}+(x_{B}-2)^{2}],$
$\frac{dx_{B}}{dt}$ $=~{}x_{A}^{2}x_{B}+x_{B}^{3}-2x_{A}x_{B}-4x_{B}^{2}+5x_{B}~{}=:f_{B}$
$=~{}x_{B}[(x_{A}-1)^{2}+(x_{B}-2)^{2}].$
It follows that the only positive root of $f_{A}=f_{B}=0$ is $(x_{A},x_{B})=(1,2)$, so this system has ACR in both species. However, we will ignore this obvious fact and instead apply the above procedure. Unless otherwise noted, all computations below can be checked using any computer algebra system, such as Macaulay2 [29] and Singular [18].
The positive-restriction ideal, $J=\langle f_{A},~{}f_{B},~{}x_{A}z_{A}^{2}-1,~{}x_{B}z_{B}^{2}-1\rangle$ in $\mathbb{Q}[x_{A},x_{B},z_{A},z_{B}]$, can be decomposed as follows:
$J=Q_{1}\cap Q_{2}\cap Q_{3}\cap Q_{4}~{},$
where
$Q_{1}$ $=\langle(x_{A}-1)^{2}+(x_{B}-2)^{2},~{}x_{A}z_{A}^{2}-1,~{}x_{B}z_{B}^{2}-1\rangle,$
$Q_{2}$ $=\langle x_{A},~{}(x_{A}-1)^{2}+(x_{B}-2)^{2},~{}x_{A}z_{A}^{2}-1,~{}x_{B}z_{B}^{2}-1\rangle,$
$Q_{3}$ $=\langle x_{B},~{}(x_{A}-1)^{2}+(x_{B}-2)^{2},~{}x_{A}z_{A}^{2}-1,~{}x_{B}z_{B}^{2}-1\rangle,\,{\rm and}$
$Q_{4}$ $=\langle x_{A},~{}x_{B},~{}x_{A}z_{A}^{2}-1,~{}x_{B}z_{B}^{2}-1\rangle~{}.$
We easily see that $V_{{\mathbb{C}}}(Q_{2})=V_{{\mathbb{C}}}(Q_{3})=V_{{\mathbb{C}}}(Q_{4})=\varnothing$. So, Lemma 4.10 implies that $V_{\mathbb{R}}({J})=V_{\mathbb{R}}(Q_{1})$. This ideal $Q_{1}$ does not contain a polynomial of the form $x_{A}-\alpha$ or $x_{B}-\alpha$ (where $\alpha>0$), so we apply Procedure 4.14 to $Q_{1}$.
It is straightforward to check that $\widetilde{Q}_{1}$ is a prime ideal in $\mathbb{R}[x_{A},x_{B},z_{A},z_{B}]$ with dimension 1. Hence, Step 1 is accomplished with $\ell=1$ and $P_{1}=Q_{1}$.
Next, for Step 2, we must check whether $\widetilde{Q}_{1}$ has a nonsingular zero. Accordingly, we consider the matrix of partial derivatives of the generators of $\widetilde{Q}_{1}$ (namely, $g_{1}:=x_{A}^{2}-2x_{A}+x_{B}^{2}-4x_{B}+5$, $g_{2}:=x_{A}z_{A}^{2}-1$, $g_{3}:=x_{B}z_{B}^{2}-1$):
$M~{}=~{}\left[\begin{array}[]{cccc}2x_{A}-2&2x_{B}-4&0&0\\ z_{A}^{2}&0&2x_{A}z_{A}&0\\ 0&z_{B}^{2}&0&2x_{B}z_{B}\\ \end{array}\right].$
A real zero of $\widetilde{Q}_{1}$ is non-singular if the matrix $M$ – when specialized at that zero – has rank 3, that is, one of the following $3\times 3$ minors of $M$ is nonzero:
$h_{1}$ $:=-4(x_{A}-1)x_{A}z_{A}z_{B}^{2}~{},$
$h_{2}$ $:=-4(x_{B}-2)x_{B}z_{A}^{2}z_{B}~{},$
$h_{3}$ $:=8(x_{A}-1)x_{A}x_{B}z_{A}z_{B}~{},$
$h_{4}$ $:=8(x_{B}-2)x_{A}x_{B}z_{A}z_{B}~{}.$
It follows that $\widetilde{Q}_{1}$ has a nonsingular zero if and only if at least one of the following ideals has nonempty real variety (for $i=1,2,3,4$):
$I_{i}:=\langle g_{1},g_{2},g_{3},h_{i}w-1\rangle\quad\subseteq\quad\mathbb{R}[x_{A},x_{B},z_{A},z_{B},w]~{},$
where $w$ is a new variable (we refer the reader to [4, Chapter 13]). Indeed, we find that $V_{\mathbb{R}}(I_{i})=\varnothing$ for all $i=1,2,3,4$, and so $\widetilde{Q}_{1}$ has no nonsingular zero.
We proceed to Step 3. We define the ideal $Q_{1}^{(1)}=\langle g_{1},g_{2},g_{3},h_{1},h_{2},h_{3},h_{4}\rangle$, whose real variety then coincides with the real variety of $J$, and has dimension 0 ($<1$). We again apply Steps 1 and 2, this time for $Q_{1}^{(1)}$, as follows. This ideal can be decomposed as $Q_{1}^{(1)}=P_{1}^{(1)}\cap P_{2}^{(1)}$, where $P_{1}^{(1)},P_{2}^{(1)}$ are the following two prime ideals of $\mathbb{R}[x_{A},x_{B},z_{A},z_{B}]$:
$P_{1}^{(1)}=\langle x_{A}-1,~{}x_{B}-2,~{}z_{A}-1,~{}2z_{B}^{2}-1\rangle,\quad P_{2}^{(1)}=\langle x_{A}-1,~{}x_{B}-2,~{}z_{A}+1,~{}2z_{B}^{2}-1\rangle.$
As $z_{B}\neq 0$ for any zero of $P_{i}^{(1)}$, it is straightforward to check that any zero of $P_{i}^{(1)}$ is non-singular, for $i=1,2$. As both $x_{A}-1$ and $x_{B}-2$ belong to $Q_{1}^{(1)}$ we conclude that the system has ACR in both species, with ACR-values $1$ and $2$, respectively.
4.4. Problems with standard approaches to deciding ACR
In this subsection, we give several sufficient conditions for ACR in terms of the steady-state ideal (Proposition 4.16). We show how this result can be useful (Example 4.17). However, we also show through examples that none of the sufficient conditions is necessary for ACR, and additionally that some other approaches to deciding ACR are also incomplete.
Proposition 4.16 (Sufficient conditions for ACR).
Let $I$ be the steady-state ideal of a mass-action system $(G,\kappa)$ with $s$ species. Let $i\in\{1,2,\dots,n\}$. Assume one of the following holds:
  1. (1)
    there exists $\alpha\in\mathbb{R}_{>0}$ such that $x_{i}-\alpha$ is in $I$,
  2. (2)
    there exists $\alpha\in\mathbb{R}_{>0}$ such that $x_{i}-\alpha$ is in the saturation $I:(x_{1}x_{2}\dots x_{n})^{\infty}$,
  3. (3)
    there exists $g\in I\cap\mathbb{R}[x_{i}]$ such that $g$ has a unique positive root.
Then $(G,\kappa)$ has ACR in species $X_{i}$.
Proof.
Condition (1) is a special case of (2) (and also (3)), so we begin with (2). Assume that $x_{i}-\alpha$ is in $I:(x_{1}x_{2}\dots x_{n})$, for some $\alpha>0$. This means that $(x_{1}x_{2}\dots x_{n})^{m}(x_{i}-\alpha)\in I$ for some $m\geq 0$. Now assume that $x^{*}$ is a positive steady state. Then $(x^{*}_{1}x^{*}_{2}\dots x^{*}_{n})^{m}(x^{*}_{i}-\alpha)=0$ and so (as $x_{j}^{*}>0$ for all $j$) we have $x^{*}_{i}=\alpha$. Hence, there is ACR in $X_{i}$ with ACR-value $\alpha$.
For (3), assume that $g\in I\cap\mathbb{R}[x_{i}]$ has a unique positive root $\alpha$. Let $x^{*}$ be a positive steady state. Then $g\in I$ implies that $g(x^{*}_{i})=0$ and so $x^{*}_{i}=\alpha$ since $\alpha$ is the only positive root of $g$. We conclude that $(G,\kappa)$ has ACR in species $X_{i}$ (with ACR-value $\alpha$). ∎
The following example shows how to use condition (2) in Proposition 4.16 to detect ACR.
Example 4.17.
Motivated by bifunctional enzymes in signal transduction networks, Joshi and Nguyen introduced the network below and showed it has ACR in $S_{3}$ [40, Table 1, line 6]:
(15) $\begin{split}S_{1}+E\overset{\kappa_{1}}{\underset{\kappa_{2}}{\rightleftarrows}}C_{1}\overset{\kappa_{3}}{\rightarrow}S_{2}+E\overset{\kappa_{4}}{\underset{\kappa_{5}}{\rightleftarrows}}C_{2}\overset{\kappa_{6}}{\rightarrow}S_{3}+E\\ S_{2}+C_{1}\overset{\kappa_{7}}{\underset{\kappa_{8}}{\rightleftarrows}}C_{3}\overset{\kappa_{9}}{\rightarrow}S_{3}+C_{1}\overset{\kappa_{10}}{\underset{\kappa_{11}}{\rightleftarrows}}C_{4}\overset{\kappa_{12}}{\rightarrow}S_{1}+C_{1}\end{split}$
The unusual ordering of the species $S_{2},S_{3},S_{1}$ in the second line of (15) is deliberate.
Joshi and Nguyen show the network (15) has ACR in $S_{3}$, by appealing to a result on bifunctional enzymes in futile cycles of a very specific form [40]. As for general-purpose results on ACR in the literature, they do not apply. For instance, the results of [68] can not be used, because the network has deficiency $2$. Similarly, the results of [75] can not be used, because the the network does not allow a ‘simple weakly reversible translation’.
We show next that Proposition 4.16 applies to this network. Let $x_{1},x_{2},\dots,x_{8}$ denote the concentrations of the species $S_{1},S_{2},S_{3},E,C_{1},C_{2},C_{3},C_{4}$, respectively. In what follows, we use symbolic rate constants, although Proposition 4.16 pertains to specialized values of rate constants, because the output we show below does not depend on the specific choice of the rate constants. The steady-state ideal $I$ is generated by the following eight polynomials:
$-\kappa_{1}x_{1}x_{4}+\kappa_{2}x_{5}+\kappa_{12}x_{8}~{},\,\kappa_{3}x_{5}+\kappa_{5}x_{6}-\kappa_{4}x_{2}x_{4}-\kappa_{7}x_{2}x_{5}+\kappa_{8}x_{7}~{},\\ \kappa_{6}x_{6}+\kappa_{9}x_{7}-\kappa_{10}x_{3}x_{5}+\kappa_{11}x_{8}~{},\,-\kappa_{1}x_{1}x_{4}+(\kappa_{2}+\kappa_{3})x_{5}-\kappa_{4}x_{2}x_{4}+(\kappa_{5}+\kappa_{6})x_{6}~{},\\ \kappa_{1}x_{1}x_{4}-(\kappa_{2}+\kappa_{3})x_{5}-\kappa_{7}x_{2}x_{5}+(\kappa_{8}+\kappa_{9})x_{7}-\kappa_{10}x_{3}x_{5}+(\kappa_{11}+\kappa_{12})x_{8}~{},\\ \kappa_{4}x_{2}x_{4}-(\kappa_{5}+\kappa_{6})x_{6}~{},\,\kappa_{7}x_{2}x_{5}-(\kappa_{8}+\kappa_{9})x_{7}~{},\,\kappa_{10}x_{3}x_{5}-(\kappa_{11}+\kappa_{12})x_{8}~{}.$
The saturation $I:(x_{1}x_{2}\dotsm x_{8})^{\infty}$ contains the polynomial $\kappa_{10}\kappa_{12}x_{3}-\kappa_{3}(\kappa_{11}+\kappa_{12})$. Thus, by Proposition 4.16(2), there is ACR in $S_{3}$ with ACR-value $\kappa_{3}(\kappa_{11}+\kappa_{12})/\kappa_{10}\kappa_{12}$.
The next examples show that the sufficient conditions for ACR in Proposition 4.16 are not necessary. (We remark that the authors – and other researchers as well [58] – once mistakenly believed condition (1) in the proposition to be necessary for ACR!) We begin by showing through the following examples that conditions (1) and (3) are not necessary for ACR.
Example 4.18.
Consider the following network:
$A+B\to 2B\quad\quad B\to A~{}.$
The ODEs are ${dx_{A}}/{dt}=-{dx_{B}}/{dt}=-\kappa_{1}x_{A}x_{B}+\kappa_{2}x_{B}=x_{B}(-\kappa_{1}x_{A}+\kappa_{2})$, so $x_{A}^{*}=\kappa_{2}/\kappa_{1}$ for all positive steady states $(x_{A}^{*},x_{B}^{*})$. Thus, there is ACR in species $A$. Nevertheless, eliminating $x_{B}$ from the steady-state ideal $I=\langle x_{B}(-\kappa_{1}x_{A}+\kappa_{2})\rangle$ yields the zero ideal: $I\cap\mathbb{R}[x_{A}]=\{0\}$. Therefore, conditions (1) and (3) of Proposition 4.16 are violated.
Example 4.19.
Consider the following network:
$A\to B\to C\quad\quad 2C\to B+C\quad\quad A+C\leftrightarrows D\to 2A~{}.$
The polynomial $\kappa_{6}\kappa_{4}x_{C}x_{D}+(-\kappa_{1}\kappa_{5}-\kappa_{1}\kappa_{6})x_{D}$ is in the steady-state ideal, so $x_{C}^{*}=(-\kappa_{1}\kappa_{5}-\kappa_{1}\kappa_{6})/(\kappa_{6}\kappa_{4})$ for all positive steady states $(x_{A}^{*},x_{B}^{*},x_{C}^{*},x_{D}^{*})$. Thus, there is ACR in species $C$. Nevertheless, eliminating $x_{D}$ from the steady-state ideal yields the zero ideal ($I\cap\mathbb{R}[x_{C}]=\{0\}$). Therefore, conditions (1) and (3) of Proposition 4.16 are violated.
From Examples 4.184.19, we might propose to allow ourselves to divide by $x_{B}$ and $x_{D}$, respectively, as Dexter et al. did [19]; mathematically, this means to saturate the steady-state ideal. Doing so would allow us to find, for the systems in Example 4.184.19, a polynomial of the desired form (namely, $-\kappa_{1}x_{A}+\kappa_{2}$ and $\kappa_{6}\kappa_{4}x_{C}+(-\kappa_{1}\kappa_{5}-\kappa_{1}\kappa_{6})$, respectively). In other words, for these two examples, condition (2) of Proposition 4.16 holds. Nonetheless, the next example shows that condition (2) fails for some other networks with ACR.
Example 4.20 (Example 2.9, continued).
Recall that the network $G=\{B\overset{\kappa_{1}}{\to}A,~{}2A+B\overset{\kappa_{2}}{\to}A+2B\}$ has ACR in species $A$. We consider fixed values $\kappa_{i}>0$ for the rate constants (the subsequent analysis does not depend on their precise values). The ODEs are ${dx_{A}}/{dt}=-{dx_{B}}/{dt}=x_{B}(\kappa_{1}-\kappa_{2}x_{A}^{2})$. After saturating the steady-state ideal by $\langle x_{A}x_{B}\rangle$, we obtain $\langle\kappa_{1}-\kappa_{2}x_{A}^{2}\rangle$. However, this ideal does not contain $x_{A}-\sqrt{\kappa_{1}/\kappa_{2}}$, nor any other polynomial of the form $x_{A}-\alpha$ where $\alpha>0$. In other words, condition (2) of Proposition 4.16 is violated.
From Example 4.20, we might propose to factor $\kappa_{1}-\kappa_{2}x_{A}^{2}$; indeed, $x_{A}-\sqrt{\kappa_{1}/\kappa_{2}}$ is a zero-divisor of the steady-state ideal. One might therefore ask whether having a unique such zero-divisor is necessary or sufficient for ACR. In other words, we consider the following condition (where the notation is as in Proposition 4.16):
(16) there exists a unique $\textrm{there exists a unique }\alpha\in\mathbb{R}_{>0}\textrm{ such that }x_{i}-\alpha\textrm{ is in }I\textrm{ or }x_{i}-\alpha\textrm{ is azero-divisor of }I.$
However, the next examples show that condition (16) is neither necessary nor sufficient for ACR.
Example 4.21.
Recall the system from Example 4.15, which generates the following ODEs:
$\frac{dx_{A}}{dt}$ $=x_{A}[(x_{A}-1)^{2}+(x_{B}-2)^{2}]$
$\frac{dx_{B}}{dt}$ $=x_{B}[(x_{A}-1)^{2}+(x_{B}-2)^{2}].$
Recall that the only positive steady state is $(1,2)$, so this system $(G,\kappa)$ has ACR in both species. However, $x_{A}-1$ is not a zero divisor of the steady state ideal; this is easily checked computationally (as explained in Remark 2.26) or alternatively a direct proof can be given. We conclude that condition (16) is not necessary for ACR.
Example 4.22.
Consider the following network:
[…]
The resulting mass-action ODEs admit the following factorization:
$\frac{dx_{A}}{dt}~{}$ $=~{}(x_{A}-1)(x_{A}-x_{B})$
$\frac{dx_{B}}{dt}~{}$ $=~{}(x_{A}-x_{B})(x_{A}+x_{B})~{}.$
The polynomial $(x_{A}-1)$ is the unique zero-divisor of the steady-state ideal $I$ of the form $x_{A}-\alpha$ (see Lemma 4.23 below). Nonetheless, there is no ACR. Indeed, the set of steady states is given by a line, $x_{A}=x_{B}$, and so $(1,1)$ and $(2,2)$ are positive steady states that differ in both coordinates. Hence, condition (16) is not sufficient for ACR.
Lemma 4.23.
The ideal $I=\langle(x-1)(x-y),~{}(x-y)(x+y)\rangle\subseteq{\mathbb{Q}}[x,y]$ has a unique zero-divisor of the form $x-\alpha$, with $\alpha\in\mathbb{Q}$, namely, $x-1$.
Proof.
It is straightforward to check that $(x-1)$ is a zero-divisor of $I$, so we need only show it is the unique such zero-divisor. To this end, assume that $(x-\alpha)$ is a zero-divisor. We must show that $\alpha=1$.
The reduced Gröbner basis of $I$ with respect to the lexicographic order $y>x$ consists of the following two polynomials:
$g_{1}~{}$ $=~{}yx-y-x^{2}+x\quad\quad{\rm and}\quad\quad g_{2}~{}=~{}y^{2}-x^{2}~{}.$
It follows that the leading ideal of $I$ is $\langle yx,y^{2}\rangle$, and so contains all monic monomials except $1,y,x,x^{2},x^{3},\dots$.
As $(x-\alpha)$ is a zero-divisor, there exists $h(x,y)\in\mathbb{Q}[x,y]\smallsetminus I$ such that $f=(x-\alpha)h(x,y)\in I$. By considering the remainder after division by the Gröbner basis, we may assume that $h$ has the form $h(x,y)=a_{1}y+c_{0}+c_{1}x+c_{2}x^{2}+\dots+c_{n}x^{n}$, where $a_{1},c_{0},\dots,c_{n}\in\mathbb{Q}$. Next, $a_{1}\neq 0$, as otherwise the leading term of $f=(x-\alpha)h(x,y)$ would not be in the leading ideal of $I$.
Consider $f-a_{1}g_{1}$, which is in $I$. It is straightforward to check that if $\alpha\neq 1$, then the leading term of $f-a_{1}g_{1}$ is $a_{1}(1-\alpha)y$, which is not in the leading ideal of $I$, and so is a contradiction. We conclude that, as desired, $\alpha=1$. ∎
Example 4.22 required some effort (via Lemma 4.23) to show that condition (16) is not sufficient for ACR. In the following example, it is easier to show the same non-sufficiency, but involves ODEs of degree $3$ (as opposed to those of degree $2$ in Example 4.22).
Example 4.24.
Consider the following ODEs, which admit a factorization:
$\frac{dx_{A}}{dt}~{}$ $=~{}(x_{A}-1)(x_{A}-x_{B})$
$\frac{dx_{B}}{dt}~{}$ $=~{}x_{B}(x_{A}-1)(x_{A}-x_{B})~{}.$
These are mass-action ODEs (by Remark 2.10). In this case, the steady-state ideal $I$ is (principally) generated by $(x_{A}-1)(x_{A}-x_{B})$, and so $x_{A}-1$ is the unique zero-divisor of $I$ of the form $x_{A}-\alpha$. Nonetheless, there is no ACR. Indeed, the set of steady states is the union of two lines – namely, $x_{A}=1$ and $x_{A}=x_{B}$ – and so $(1,1)$ and $(2,2)$ are positive steady states that differ in both coordinates. Hence, condition (16) is not sufficient for ACR.
The prior three examples pertain to networks with complexes that are beyond bimolecular (like $3A+B$). Indeed, Examples 4.21 and 4.24 involve degree-3 polynomials. On the other hand, Example 4.22 features degree-2 polynomials only (the trimolecular complex, $3A+B$, is not a reactant). A natural question is whether there are examples with lower molecularity:
Question 4.25.
For networks that are at-most-bimolecular, is condition (16) necessary for ACR? Is it sufficient?
Biochemical networks arising in applications are typically at-most-bimolecular, so answers to Question 4.25 could be of great practical importance.
A final approach to deciding ACR, is to try sampling steady states from a few compatibility classes, but this too might not suffice.
Example 4.26.
Consider the following network:
[…]
The ODEs factor as shown here:
$\frac{dx_{A}}{dt}~{}$ $=~{}x_{B}(x_{A}-1)(x_{A}-2)$
$\frac{dx_{B}}{dt}~{}$ $=~{}-x_{B}(x_{A}-1)(x_{A}-2)~{}.$
Hence, the set of steady states has two components, arising from $x_{A}=1$ and $x_{A}=2$ (see Figure 1). If we sampled a few stoichiometric compatibility classes, and checked whether the steady-state value of $x_{A}$ is always the same, we could get unlucky – and conclude erroneously that $A$ had ACR with ACR-value $1$.
000.50.50.50.511111.51.51.51.522222.52.52.52.5333300111122223333
Figure 1. For the system arising from (17), the positive steady-state locus consists of two components (the vertical, red lines), so sampling from some of the compatibility classes (blue dashed lines) could erroneously lead to concluding that there is ACR.
As before, Example 4.26 involves a degree-3 polynomial. So, we again would like to know whether there are similar examples involving lower-degree polynomials.
We end this section with an example of a network arising in biology that exhibits ACR.
Example 4.27 (Shinar and Feinberg network).
The following network was analyzed (and shown to have ACR) by Shinar and Feinberg [68, Figure 2(B)] and has been studied by many others [2, 20, 25, 42, 64]:
(18) $\begin{split}X\overset{\kappa_{1}}{\underset{\kappa_{2}}{\rightleftarrows}}XT\overset{\kappa_{3}}{\rightarrow}X_{p}\quad\quad X_{p}+Y\overset{\kappa_{4}}{\underset{\kappa_{5}}{\rightleftarrows}}X_{p}Y\overset{\kappa_{6}}{\rightarrow}X+Y_{p}\quad\quad XT+Y_{p}\overset{\kappa_{7}}{\underset{\kappa_{8}}{\rightleftarrows}}XTY_{p}\overset{\kappa_{9}}{\rightarrow}XT+Y\end{split}$
We denote by $x_{1},x_{2},\dots,x_{7}$ the concentrations of the species as follows:
$x_{X}=x_{1},\;x_{XT}=x_{2},\;x_{X_{p}}=x_{3},\;x_{Y}=x_{4},\;x_{Y_{p}}=x_{5},\;x_{X_{p}Y}=x_{6},\;x_{XTY_{p}}=x_{7}~{}.$
The steady state ideal $I$ is generated by $-\kappa_{1}x_{1}+\kappa_{2}x_{2}+\kappa_{6}x_{6},~{}\kappa_{3}x_{2}-\kappa_{4}x_{3}x_{4}+\kappa_{5}x_{6},~{}\kappa_{6}x_{6}-\kappa_{7}x_{2}x_{5}+\kappa_{8}x_{7},~{}\kappa_{4}x_{3}x_{4}-(\kappa_{5}+\kappa_{6})x_{6},~{}\kappa_{7}x_{2}x_{5}-(\kappa_{8}+\kappa_{9})x_{7}$. The reduced Gröbner basis of $I$ with respect to the lexicographic order $x_{1}>x_{2}>x_{4}>x_{6}>x_{7}>x_{5}$ is
${\mathcal{G}}~{}$ =$=~{}\{\kappa_{7}\kappa_{9}x_{5}x_{7}-(\kappa_{3}\kappa_{8}+\kappa_{3}\kappa_{9})x_{7},~{}\kappa_{6}x_{6}-\kappa_{9}x_{7},~{}\kappa_{4}x_{3}x_{4}-(\kappa_{5}+\kappa_{6})x_{6},~{}$
$\quad\quad\quad\quad\kappa_{3}x_{2}-\kappa_{6}x_{6},~{}\kappa_{1}x_{1}-\kappa_{2}x_{2}-\kappa_{6}x_{6}\}~{}.$
Here, ACR in $X_{5}$ can be seen from the first Gröbner basis element, $g_{1}:=\kappa_{7}\kappa_{9}x_{5}x_{7}-(\kappa_{3}\kappa_{8}+\kappa_{3}\kappa_{9})x_{7}$. Indeed, $g_{1}/x_{7}=\kappa_{7}\kappa_{9}x_{5}-(\kappa_{3}\kappa_{8}+\kappa_{3}\kappa_{9})$ is a linear polynomial in $x_{5}$ and so, for all choices of rate constants, this network exhibits ACR in the fifth species, namely, $Y_{p}$.
Remark 4.28.
One message of this subsection is that, in general, ACR is not readily detected from the steady-state ideal. If, however, we allow for a more general form of ACR (complex-number ACR), then detection from the steady-state ideal is possible. See Appendix B.
5. Numerical methods for detecting or precluding ACR
In Section 2.4, we gave necessary and sufficient conditions for ACR in a reaction system $(G,\kappa)$ in terms of the positive steady-state locus $>0]{I(G,\kappa)}$. In particular, Proposition 2.15 states that $(G,\kappa)$ has ACR in $X_{i}$ if and only if $x_{i}-\alpha$ is in $>0]{I(G,\kappa)}$ for some $\alpha>0$. As discussed throughout, the fundamental challenge to detecting ACR is that it requires an understanding of the ideal $>0]{I(G,\kappa)}$. While a few (restricted) techniques exist to fully understand $>0]{I(G,\kappa)}$, especially when working over $\mathbb{R}$ rather than $\mathbb{Q}$, we can gain some knowledge using techniques from numerical algebraic geometry.
Numerical algebraic geometry is a field concerned with computing solutions to a system of polynomial equations, often using algorithms based on polynomial homotopy continuation. The fundamentals of polynomial homotopy continuation can be found in [7]. Algorithms in numerical algebraic geometry not only focus on finding isolated solutions to systems, but also positive-dimension solution sets, e.g. curves and surfaces. In essence, these algorithms give a geometric way to decompose a complex variety, and with care, can yield insight into the decomposition of a real variety.
In this section, we start by providing background on real algebraic sets (Section 5.1) and describing some tools and objects from numerical algebraic geometry, in particular, numerical irreducible decompositions and witness sets (Section 5.2). We then use these tools to give an algorithm for detecting ACR (Procedure 5.7) and an algorithm for precluding ACR (Procedure 5.10). The first uses numerical irreducible decompositions and witness sets to see whether any component of $V_{\mathbb{C}}(I(G,\kappa))$ lies in a hyperplane of the form $\{x_{i}-\alpha\}$ (Section 5.3). The second uses tools for numerically solving polynomial equations to find a dense set of points on each component of $V_{\mathbb{R}}(I(G,\kappa))$ that intersects a specified region of the positive orthant (Section 5.4).
Remark 5.1.
In this section, we work with the steady-state ideal $I(G,\kappa)$, rather than the positive-restriction ideal $J(G,\kappa)$ (from Definition 4.3). This choice allows us to keep the number of equations and variables to half of those in the positive-restriction ideal. Nevertheless, both of the algorithms we propose in this section can be used with $J(G,\kappa)$.
5.1. Real algebraic sets
We start this section by introducing real algebraic sets and restating Proposition 2.15 from a geometric point-of-view (see Proposition 5.2).
A real algebraic set, or real algebraic variety, is the intersection of a complex variety with $\mathbb{R}^{n}$. For instance, consider the real variety
$V_{\mathbb{R}}(\sqrt[>0]{I(G,\kappa)})~{}=~{}\{x\in\mathbb{R}^{n}\ |\ h(x)=0\text{ for all }h\in\sqrt[>0]{I(G,\kappa)}\}~{}.$
We can rewrite this variety as the following interesection:
$V_{\mathbb{R}}(\sqrt[>0]{I(G,\kappa)})~{}=~{}\mathbb{R}^{n}\cap V_{\mathbb{C}}(\sqrt[>0]{I(G,\kappa)})~{},$
where
$V_{\mathbb{C}}(\sqrt[>0]{I(G,\kappa)})~{}=~{}\{x\in\mathbb{C}^{n}\ |\ h(x)=0\text{ for all }h\in\sqrt[>0]{I(G,\kappa)}\}~{}.$
Proposition 5.2.
A reaction system $(G,\kappa)$ has ACR in $X_{i}$ if and only if there exists $\alpha>0$ such that $V_{\mathbb{R}}(\sqrt[>0]{I(G,\kappa)})$ lies in the hyperplane $\{x_{i}-\alpha=0\}$.
Proof.
This result follows immediately from Proposition 2.15. ∎
As we explain in the next subsection, every complex variety can be written as the union of irreducible components. Real algebraic sets can also be decomposed. Indeed, a real algebraic set $V_{\mathbb{R}}$ can be decomposed into the union of finitely many path-connected sets $C_{1},C_{2},\ldots,C_{\ell}$ where, for each $i$, both $C_{i}$ and $V_{\mathbb{R}}\smallsetminus C_{i}$ are closed in the Euclidean topology. We call each such set $C_{i}$ a real connected component of $V_{\mathbb{R}}$ (more details on real connected components can be found in [9]). The following result pertains to connected components of $V_{\mathbb{R}}(\sqrt[>0]{I(G,\kappa)})$, and it follows immediately from Proposition 5.2.
Corollary 5.3.
A reaction system $(G,\kappa)$ has ACR in $X_{i}$ if and only if there exists $\alpha>0$ such that every real connected component of $V_{\mathbb{R}}(\sqrt[>0]{I(G,\kappa)})$ lies in the hyperplane $\{x_{i}-\alpha=0\}$.
Recall from (8) the containments $I(G,\kappa)~{}\subseteq~{}\sqrt{I(G,\kappa)}~{}\subseteq~{}\sqrt[>0]{I(G,\kappa)}$. It follows that
$V_{\mathbb{C}}(\sqrt[>0]{I(G,\kappa)})\cap\mathbb{R}^{n}~{}\subseteq~{}V_{\mathbb{C}}(\sqrt{I(G,\kappa)})\cap\mathbb{R}^{n}~{}\subseteq~{}V_{\mathbb{C}}(I(G,\kappa))\cap\mathbb{R}^{n},$
and so each real connected component of $V_{\mathbb{R}}(\sqrt[>0]{I(G,\kappa)})$ is contained in a real connected component of $V_{\mathbb{R}}(I(G,\kappa))$. This fact, together with Corollary 5.3, yields the following sufficient condition to detect ACR.
Proposition 5.4.
If there exists $\alpha>0$ such that every real connected component of $V_{\mathbb{R}}(I(G,\kappa))$ lies either in some coordinate hyperplane $\{x_{j}=0\}$ or in the hyperplane $\{x_{i}-\alpha=0\}$, then the reaction system $(G,\kappa)$ has ACR in $X_{i}$.
The condition in Proposition 5.4 is sufficient, but not necessary for ACR. For example, if $I(G,\kappa)=\langle x_{A}x_{B}(x_{A}-1)(x_{B}+2)\rangle$, then $V_{\mathbb{R}}(I(G,\kappa))$ has four real components: the real hyperplanes $\{x_{A}=0\}$, $\{x_{B}=0\}$, $\{x_{A}-1=0\}$, and $\{x_{B}+2=0\}$. By examining the last component, $\{x_{B}+2=0\}$, we see that the reaction system $(G,\kappa)$ does not satisfy the hypothesis of Proposition 5.4, despite having ACR in species A.
In this section, we give two numerical techniques to check for ACR. Procedure 5.10 – for precluding ACR – relies on Proposition 5.4, while Procedure 5.7 – for detecting ACR – relies on numerical irreducible decompositions, which we define in the next subsection.
5.2. Numerical irreducible decomposition and witness sets
Let $V_{{\mathbb{C}}}$ be a complex variety. The pure $j$-dimensional component of $V_{{\mathbb{C}}}$ is the union of the irreducible components of dimension $j$. Thus, the variety $V_{{\mathbb{C}}}$ has an irreducible decomposition of the form
(19) $V_{{\mathbb{C}}}~{}=~{}\bigcup_{j=0}^{\text{dim}V_{{\mathbb{C}}}}V_{j}~{}=~{}\bigcup_{j=0}^{\text{dim}V_{{\mathbb{C}}}}\bigcup_{l=1}^{k_{j}}V_{j,l}~{},$
where $V_{j}$ is the pure $j$-dimensional component of $V_{{\mathbb{C}}}$, and each $V_{j,l}$ is a $j$-dimensional irreducible component.
Irreducible decompositions (19) can be found using symbolic or numerical methods. Symbolic methods, which are more costly, find irreducible decompositions algebraically, and the output is a set of ideals described by generators. In contrast, methods using numerical algebraic geometry find numerical irreducible decompositions by combining tools from algebraic geometry and numerical analysis (e.g., via a polynomial homotopy continuation), and the output is a witness set.
Here we briefly describe numerical irreducible decompositions (see [7] for more information). At a high level, a numerical irreducible decomposition is a collection of points obtained from intersecting each complex component of the variety with a generic linear space. In particular, let $V_{j,l}$ be a $j$-dimensional irreducible component of $V_{{\mathbb{C}}}$ with $d_{j,l}=\deg~{}V_{j,l}$, then for a fixed $j$-codimensional generic linear space $H_{j,l}\subseteq{\mathbb{C}}^{n}$, we have that the intersection $V_{j,l}\cap H_{j,l}$ consists of $d_{j,l}$ points.
For ease of reference, we associate to each $H_{j,l}$ a linear system $L_{j,l}$. Specifically, $L_{j,l}:{\mathbb{C}}^{n}\rightarrow{\mathbb{C}}^{j}$ is a system of linear polynomials such that $V_{{\mathbb{C}}}(L_{j,l})=H_{j,l}$. The set $V_{j,l}\cap V_{{\mathbb{C}}}(L_{j,l})=V_{j,l}\cap H_{j,l}$ is called a witness point set for $V_{j,l}$. The triple $\mathcal{W}_{j,l}=\{I(V_{{\mathbb{C}}}),L_{j,l},V_{j,l}\cap V_{{\mathbb{C}}}(L_{j,l})\}$ is a witness set for $V_{j,l}$. A numerical irreducible decomposition of $V_{{\mathbb{C}}}$ is of the form
(20) $\bigcup_{j=0}^{\text{dim}V_{{\mathbb{C}}}}\bigcup_{l=1}^{k_{j}}\mathcal{W}_{j,l}~{},$
where $\mathcal{W}_{i,l}$ is a witness set for a distinct $j$-dimensional irreducible component of $V_{{\mathbb{C}}}$. We note that the union in (20) should be considered as a formal union. Also, by construction, each witness set includes information about the dimension and degree of the corresponding component. In the next subsection, we show how numerical irreducible decompositions can be used to detect ACR.
Remark 5.5.
Numerical irreducible decompositions can be computed using software such as Bertini [6], PHCpack [77], and HomotopyContinuation.jl [10].
5.3. ACR via numerical irreducible decomposition
Consider the complex variety of the steady-state ideal of a mass-action system $(G,\kappa)$ with $n$ species:
$V_{{\mathbb{C}}}^{G,\kappa}~{}=~{}V(I(G,\kappa))~{}=~{}\{x\in{\mathbb{C}}^{n}\mid f_{\kappa}(x)_{1}=f_{\kappa}(x)_{1}=\dots f_{\kappa}(x)_{n}=0\}~{}.$
Recall that a numerical irreducible decomposition of $V_{{\mathbb{C}}}^{G,\kappa}$ will return a witness set collection, where each witness set has the form $\mathcal{W}_{j,l}=\{I(V_{{\mathbb{C}}}^{G,\kappa}),L_{j,l},V_{j,l}\cap V_{{\mathbb{C}}}(L_{j,l})\}$ (and $V_{j,l}\cap V_{{\mathbb{C}}}(L_{j,l})$ is a finite set of points). Furthermore, if we compute a numerical irreducible decomposition of $V_{{\mathbb{C}}}^{G,\kappa}$ using a generic linear space that is different from $V_{{\mathbb{C}}}(L_{j,l})$, then the set of points in $V_{j,l}\cap V_{{\mathbb{C}}}(L_{j,l})$ will change, unless $j=0$ (that is, unless $V_{j,l}$ is a 0-dimensional component). We state this observation formally, as follows.
Proposition 5.6.
Let $\alpha\in\mathbb{C}$. Let $V_{\mathbb{C}}$ be an irreducible complex variety of dimension $d$. Let $H_{1}$ and $H_{2}$ be two distinct generic codimension-$d$ linear spaces. If there exist $p\in V_{\mathbb{C}}\cap H_{1}$ and $q\in V_{\mathbb{C}}\cap H_{2}$ such that $p_{i}=q_{i}=\alpha$ for some $i$, then $V_{\mathbb{C}}$ lies in the hyperplane $\{x_{i}-\alpha=0\}$.
Proof.
Follows from the genericity of $H_{1}$ and $H_{2}$. ∎
It follows from Proposition 5.6 that we can detect ACR by computing two (distinct) numerical irreducible decompositions and then comparing results. This observation underlies the next procedure. As the procedure requires checking equality up to an estimation error $\delta$, we refer to the ACR detected by the procedure as Numerical ACR within $\delta$. Note though, due to estimation error, it is possible that the procedure falsely detects ACR.
Procedure 5.7.
Detecting Numerical ACR by numerical irreducible decomposition
  • Input.
    A reaction system $(G,\kappa^{*})$ with ideal $I(G,\kappa^{*})$ generated by steady-state equations $f_{\kappa^{*}}(x)~{}=~{}(f_{\kappa^{*}}(x)_{1},f_{\kappa^{*}}(x)_{2},\dots,f_{\kappa^{*}}(x)_{n})$. Numerical tolerance $\delta>0$.
  • Output.
    “Numerical ACR within $\delta$” or “Inconclusive.”
  • Step 1.
    Compute two distinct numerical irreducible decompositions of $V_{{\mathbb{C}}}^{G,\kappa}$ and store witness sets in two lists, $\mathcal{W}$ and $\mathcal{W}^{\prime}$.
  • Step 2.
    Remove witness sets that represent boundary components from $\mathcal{W}$ and $\mathcal{W}^{\prime}$. In particular, for each witness set $\mathcal{W}_{j,l}=\{I(V_{{\mathbb{C}}}),L_{j,l},V_{j,l}\cap V_{{\mathbb{C}}}(L_{j,l})\}$, if there exists $1\leq i\leq n$ such that for all points $(x_{1},x_{2},\ldots,x_{n})\in V_{j,l}\cap V_{{\mathbb{C}}}(L_{j,l})$ it is the case that $|x_{i}|<\delta$, then update $\mathcal{W}$ by removing $\mathcal{W}_{j,l}$. Repeat the process for $\mathcal{W}^{\prime}$.
  • Step 3.
    If there exists a coordinate $x_{i}$ such that for all the points in $V_{j,l}\cap V_{{\mathbb{C}}}(L_{j,l})$, ranging over all $\mathcal{W}_{j,l}\in\mathcal{W}$, and for all the points in $V_{j,l}\cap V_{{\mathbb{C}}}(L^{\prime}_{j,l})$, ranging over all $\mathcal{W}^{\prime}_{j,l}\in\mathcal{W}^{\prime}$, the absolute value of pairwise differences of the $i$th coordinates is less than $\delta$, then return “Numerical ACR in $X_{i}$ within $\delta$.”
  • Step 4.
    Otherwise, return “Inconclusive.”
Procedure 5.7 returns “Inconclusive” in the case where there are two points in the returned witness sets that differ at every coordinate. In this case, since we are working over complex space, we cannot preclude ACR because we do not know how each component intersects the positive real orthant.
Proof of correctness of Procedure 5.7.
This follows from Propositions 5.2 and 5.6. ∎
Example 5.8.
(Lee et al. Wnt model) Here we use Procedure 5.7 to test whether a model of Wnt signalling has ACR. The network is displayed in the input below.
We begin with some biological context and a description of the network. Signalling pathways, such as Wnt/$\beta$-catenin, are crucial to normal functioning of cells, with many diseases such as cancer caused by dysfunction of the pathway. Lee et al. [48] introduced the first quantitative model of Wnt/$\beta$-catenin based on data from Xenopus extracts. The model describes known interactions between core components of the canonical pathway where species can become active or inactive, denoted by a subscript $a$ or $i$, respectively.
[…]
As far as we are aware, our analysis is the first testing of ACR in the Lee et al. Wnt model. To do so, we use Macaulay2 [29] calling the ReactionNetworks.m2 package [33], the NumericalAlgebraicGeometry.m2 package [49], and the following code. In the code, we substitute the reaction rates with random rational positive values. We use $\delta=10^{-8}$ since that is the default error tolerance in NumericalAlgebraicGeometry.m2.
RN=  reactionNetwork({"Di<-->Da", "Ya<-->Yi", "Da+Yi-->Da + G + CNA",
 "G + CNA --> Yi", "Yi --> G + CNA", "A+N<-->CNA", "Ya+X<-->CXY",
 "CXY-->CYXp", "CYXp --> Xp + Ya",Ψ"Xp-->0", "X<-->0", "0<-->N",
 "X+T<-->CXT", "X+A<-->CXA"
         }, NullSymbol => "0");
R = createRing(RN,QQ);
I = ideal( subRandomReactionRates RN);
S=QQ[RN.ConcentrationRates];
J = sub(I, S);
NV1 = numericalIrreducibleDecomposition(J);
Components1=components NV1;
This returns to us the following numerical variety with a single witness set representing a single irreducible component of dimension 4 and degree 7:
i1: peek NV1
o1= NumericalVariety{4 => {(dim=4,deg=7)}}
Taking a peek at the witness point of the component gives us
i1: (points Components1_0)_0
o1 =  {-.257725+.843632*ii, -1.23708+4.04943*ii,
-.351557-.187174*ii, -2.87105-1.52859*ii, -.415284+3.13649*ii,
-1.6121-.443075*ii, -.286596-.0787689*ii, 3,
-.256615+.671232*ii, .479671-.417656*ii, 1.5989-1.39219*ii,
1.19918-1.04414*ii, -1.52228+3.79967*ii, -.493674-.456425*ii,
3.16043-4.30398*ii}
We can see that there is a $3$ in the 8th entry, which corresponds to Axin ($N$). In fact, in this example, all seven witness points contains a $3$ in the 8th entry. Here is a second witness point to inspect.
i1: (points Components1_0)_1
o1 =  {-.127239-.106091*ii,  -.610746-.509235*ii,
-.071024-.0709828*ii, -.580029-.579693*ii, .65341+.178855*ii,
-.538247-.022909*ii, -.0956883-.0040726*ii, 3,
.617796+.209846*ii, -.0644063-.130571*ii, -.214688-.435237*ii,
-.161016-.326428*ii, .547114+1.75404*ii, -.006874+.273931*ii,
-1.45653-.564897*ii}
Comparing witness points can give us a way to immediately get a sense if ACR can be detected. We can confirm the results by performing a second numerical irreducible decomposition. This amounts to choosing a new generic linear space $V_{{\mathbb{C}}}(L_{4,1})$. Again, we get a numerical variety of dimension 4 and degree 7.
i1: NV2 = numericalIrreducibleDecomposition(J);
i2: peek NV2
o2= NumericalVariety{4 => {(dim=4,deg=7)}}
Furthermore, every witness point has a 3 in the 8th entry.
i1: (points Components2_0)_1
o1 =  {-.691939+.081125*ii, -3.32131+.389401*ii,
.434359+.0253714*ii, 3.54727+.2072*ii, -.182904-.70573*ii,
.307375-3.42684*ii, .054645-.609217*ii, 3,
.201325-.0071503*ii,.194731+.0044491*ii, .649103+.014830*ii,
.486827+.0111228*ii, -3.10059+6.38549*ii, -.132244+.29891*ii,
.16613-3.07604*ii}
Since each coordinate corresponding to Axin ($N$) is within $\delta$ for all points in both witness sets, we conclude that the system has Numerical ACR in N within $\delta=10^{-8}$.
5.4. ACR via numerically computing points on each real connected component
While finding the real connected components of a real algebraic set symbolically is challenging, we can get some information about the connected components by using numerical algebraic geometry. In particular, using results from [24, 32, 67], we can find at least one point on each real connected component, and by repeating this procedure we can find multiple points on each real connected component. In fact, we can find a provably dense set of points on each real connected component [24]. This yields a numerical algorithm for precluding ACR (Procedure 5.10).
Given a reaction system $(G,\kappa)$ with ideal $I(G,\kappa)$ generated by steady-state equations $f_{\kappa}(x)=(f_{\kappa}(x)_{1},f_{\kappa}(x)_{2},\dots,f_{\kappa}(x)_{n})$, we can find a real point on every component of $V_{\mathbb{R}}(I(G,\kappa))$ using an approach by Seidenberg [67]. In that approach, we let $g=f_{\kappa}(x)_{1}^{2}+f_{\kappa}(x)_{2}^{2}+\dots+f_{\kappa}(x)_{n}^{2}$, and choose a random point $w\in\mathbb{R}^{n}$. We then solve the following system in indeterminates $x$ and $\lambda=(\lambda_{0},\lambda_{1})$:
(21) $g(x)~{}$ $=~{}0,$
$\lambda_{0}(x-w)+\lambda_{1}\nabla g~{}$ $=~{}0~{}.$
For the system (21), which is referred to as the Fritz John optimality conditions, the solutions will include the closest point on $V_{\mathbb{R}}(I(G,\kappa))$ to the point $w$. As described in [32], solutions to the Fritz John optimality conditions can be found through constructing and, in turn, tracking a homotopy. For $\beta\in\mathbb{C}^{n-d}$, the homotopy we want to consider is
(22) $H_{w,\beta}(x,\lambda,t)=\left[\begin{array}[]{c}g(x)-t\beta\\ \lambda_{0}(x-w)+\lambda_{1}\nabla g=0\end{array}\right].$
For generic choices of $w$ and $\beta$, the homotopy $H_{w,\beta}$ is a well-constructed homotopy [24, Proposition 3.2]. In practice, this means that we can find the solutions to the system in (22) by tracking a finite number of paths from the solutions $H_{w,\beta}(x,\lambda,1)$ at $t=1$ to the target solutions $H_{w,\beta}(x,\lambda,0)$ at $t=0$, which are solutions of (21). Such homotopies can be tracked using polynomial homotopy continuation solvers Bertini [6], PHCpack [77], and HomotopyContinuation.jl [10].
Repeatedly choosing pairs $(w,\beta)$, tracking the resulting homotopy (22), and projecting target solutions at $t=0$ onto the $x$-coordinates gives us a way to find multiple points on each component. Moreover, using Algorithm 4.2 in [24], we can find a dense sample of points on $V_{\mathbb{R}}(I(G,\kappa))$ in a given region, as follows.
Lemma 5.9.
There is an algorithm, which we call DenseSamplingAlgorithm, with the following input and output:
  • Input: A set of real polynomials $f=(f_{1},f_{2},\dots,f_{k})$ on $n$ variables, a box $R=[a_{1},b_{1}]\times\cdots\times[a_{n},b_{n}]$, a sampling density $\epsilon>0$, and an estimation error $0\leq\delta<\epsilon$.
  • Output: A list $A$ of points that form an $(\epsilon,\delta)$-sample of $V_{\mathbb{R}}(f)\cap R$.
Here, an $(\epsilon,\delta)$-sample means (i) every point in $A$ is within $\delta$ of a point in $V_{\mathbb{R}}(f)\cap R$, and (ii) every point in $V_{\mathbb{R}}(f)\cap R$ is within $\epsilon$ of a point in $A$.
The next procedure shows how to numerically preclude ACR using DenseSamplingAlgorithm. This method can also be used to gain a better understanding of the components of $V_{\mathbb{R}}(\sqrt[>0]{I(G,\kappa)})$.
Procedure 5.10.
Numerically precluding ACR by computing points on each real connected component
  • Input.
    A reaction system $(G,\kappa^{*})$ with ideal $I(G,\kappa^{*})$ generated by steady-state equations $f_{\kappa^{*}}(x)~{}=~{}(f_{\kappa^{*}}(x)_{1},f_{\kappa^{*}}(x)_{2},\dots,f_{\kappa^{*}}(x)_{n})$, a box $R=[a_{1},b_{1}]\times\cdots\times[a_{n},b_{n}]$ with $0<a_{i}<b_{i}$, a sampling density $\epsilon$, and an estimation error $\delta$ with $0\leq\delta<\epsilon$.
  • Output.
    “No Numerical ACR” or “Inconclusive”
  • Step 1.
    Let $Sols$ by the output of DenseSamplingAlgorithm[$f_{\kappa}(x),R,\epsilon,\delta$].
  • Step 2.
    If $\lvert Sols\rvert=0$ or $1$, or if there exists a coordinate $x_{i}$ such that for all the points in $Sols$ the absolute value of pairwise differences of the $i$th coordinates is less than $\delta$, then return “Inconclusive”.
  • Step 3.
    Else, return “No Numerical ACR”.
Proof of correctness.
The correctness of the procedure follows from [24], Corollary 5.3, and the fact that the box $R$ is contained in the positive orthant. ∎
6. Detecting zero-divisor ACR
In earlier sections, we presented several procedures that, in some cases, detect ACR. We present a final such procedure in this section, which works well for a type of ACR we call “zero-divisor ACR”. We motivate and define this concept in Section 6.1, and then present our procedure (Algorithm 1) – and examples of its usage – in Section 6.2. The ideas behind this algorithm come from the theory of Gröbner bases for ideals involving parameters, which we describe in Section 6.3.
6.1. Gröbner bases, elimination orders, and zero-divisor ACR
Fix $i\in\{1,2,\dots,n\}$, and call $\hat{\mathbf{x}}_{i}=(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n})$. An elimination order for $\hat{\mathbf{x}}_{i}$ on ${\mathbb{Q}}[x_{1},x_{2},\dots,x_{n}]$ is a monomial order such that every polynomial with leading monomial in ${\mathbb{Q}}[x_{i}]$ belongs to ${\mathbb{Q}}[x_{i}]$. One instance of such an order is a lexicographic order with $x_{i}$ smaller than the rest of the variables. Another example is a product order where
$\hat{\mathbf{x}}_{i}^{\alpha}x_{i}^{a}\succ\hat{\mathbf{x}}_{i}^{\beta}x_{i}^{b}\;\quad\Longleftrightarrow\quad\;\hat{\mathbf{x}}_{i}^{\alpha}\succ\hat{\mathbf{x}}_{i}^{\beta},\text{ or }\hat{\mathbf{x}}_{i}^{\alpha}=\hat{\mathbf{x}}_{i}^{\beta}\text{ and }a>b.$
Given an elimination order $\succ$ for $\hat{\mathbf{x}}_{i}$ on ${\mathbb{Q}}[x_{1},x_{2},\dots,x_{n}]$, let $\succ_{\hat{\mathbf{x}}_{i}}$ denote the order on the ring ${\mathbb{Q}}[x_{i}][\hat{\mathbf{x}}_{i}]$ obtained by restricting the order $\succ$ to the monomials in ${\mathbb{Q}}[\hat{\mathbf{x}}_{i}]$. We denote the resulting leading coefficient of a polynomial $f\in\mathbb{Q}[x_{i}][\hat{\mathbf{x}}_{i}]$ as follows:
$\mathrm{lc}_{\hat{\mathbf{x}}_{i}}(f)~{}\in~{}\mathbb{Q}[x_{i}]~{}.$
One fact that we use below is the well-known Elimination Theorem. This theorem states that for any elimination order for $\hat{\mathbf{x}}_{i}$ on ${\mathbb{Q}}[x_{1},x_{2},\dots,x_{n}]$, if we compute a Gröbner basis $\mathcal{G}$ of an ideal $I\subseteq{\mathbb{Q}}[x_{1},x_{2},\dots,x_{n}]$, then $I\cap{\mathbb{Q}}[x_{i}]\neq\{0\}$ if and only if $\mathcal{G}\cap{\mathbb{Q}}[x_{i}]\neq\varnothing$, and, in this case, $I\cap{\mathbb{Q}}[x_{i}]=\langle\mathcal{G}\cap{\mathbb{Q}}[x_{i}]\rangle$ (see, for instance, [31, Theorem 4.8]).
Our aim is to use Gröbner bases to propose candidates for ACR species and their corresponding ACR-values. We motivate this approach through the following example.
Example 6.1.
[…]
Consider the lexicographic order $x_{B}>x_{C}>x_{D}>x_{A}$. The reduced Gröbner basis of $I$ with respect to this order is
(23) ${\mathcal{G}}=\{x_{A}^{2}x_{D}-4x_{A}x_{D}+3x_{D},~{}x_{A}^{2}x_{C}-3x_{A}x_{C}+2x_{C},~{}x_{A}x_{C}x_{D}-x_{C}x_{D},~{}x_{B}-x_{C}\}~{}.$
By the Elimination Theorem, we see that $I\cap{\mathbb{Q}}[x_{A}]=\{0\}$, since ${\mathcal{G}}\cap{\mathbb{Q}}[x_{A}]=\varnothing$. The leading coefficients with respect to $\hat{\mathbf{x}}_{A}=(x_{B},x_{C},x_{D})$ are, respectively:
$x_{A}^{2}-4x_{A}+3,\quad x_{A}^{2}-3x_{A}+2,\quad x_{A}-1,\quad 1~{}.$
From the third element of the Gröbner basis $\mathcal{G}$, namely, $x_{A}x_{C}x_{D}-x_{C}x_{D}=x_{C}x_{D}(x_{A}-1)$, we see that neither $x_{C}x_{D}$ nor $x_{A}-1$ is in $I$, and so $x_{A}-1$ is a zero divisor of $I$ (recall Definition 2.24). Moreover, it is easy to deduce from this polynomial that, in every positive steady state, the value of $x_{A}$ must be $1$. That is, there is ACR in $X_{A}$ with ACR-value $1$.
Example 6.1 inspires the next definition and the algorithm in the next subsection.
Definition 6.2.
Assume $\alpha>0$. A mass-action system $(G,\kappa^{*})$ has zero-divisor ACR in $X_{i}$ with ACR-value $\alpha$ if $(G,\kappa^{*})$ has ACR in $X_{i}$ with ACR-value $\alpha$ and either (1) $x_{i}-\alpha$ is in the steady-state ideal or (2) $x_{i}-\alpha$ is a zero-divisor of the steady-state ideal.
Example 6.3 (Example 4.27, continued).
We return to the Shinar and Feinberg network (18). From the Gröbner basis shown earlier, we see that the linear polynomial $g_{1}/x_{7}=\kappa_{7}\kappa_{9}x_{5}-(\kappa_{3}\kappa_{8}+\kappa_{3}\kappa_{9})$ is a zero divisor of $I$, as $g_{1}/x_{7}\notin I$, $x_{7}\notin I$, and $g_{1}\in I$. Therefore, for all choices of rate constants, this network exhibits zero-divisor ACR (in the fifth species, namely, $Y_{p}$).
Remark 6.4.
The concept of zero-divisor ACR is closely related to condition (16), which we saw earlier is neither necessary nor sufficient for ACR. In particular, we saw a mass-action system with ACR, but not zero-divisor ACR (Example 4.21).
Remark 6.5.
We saw earlier a situation where we have zero-divisor ACR, and saturating the steady-state ideal was not enough to make $x_{i}-\alpha$, where $\alpha$ is the ACR-value, appear in the ideal (Example 4.20).
Remark 6.6 (Zero-divisor ACR detection through Gröbner bases).
For many mass-action systems (or even networks), zero-divisor ACR can be detected from a reduced Gröbner basis $\mathcal{G}$ under a suitable elimination order for $\hat{\mathbf{x}}_{i}$. More precisely, in many examples, when there is zero-divisor ACR in some species $X_{i}$ with ACR-value $\alpha$, the following condition holds:
(24) there exists $\textrm{there exists }\alpha\in\mathbb{R}_{>0}\textrm{ such that }x_{i}-\alpha\textrm{ divides some element of }\mathcal{G}~{}.$
Such a system was shown in Example 6.1. As for a network, recall, in Examples 4.27 and 6.3, that $\kappa_{7}\kappa_{9}x_{5}-(\kappa_{3}\kappa_{8}+\kappa_{3}\kappa_{9})$ divides the first element of the reduced Gröbner basis of the steady-state ideal $I$ with respect to the lexicographic order $x_{1}>x_{2}>x_{4}>x_{6}>x_{7}>x_{5}$. Thus, in this example, ACR in $X_{5}$ can be detected from (factors of) the Gröbner basis, as in (24).
In light of Remark 6.6, it is natural to ask whether condition (24) is necessary for zero-divisor ACR. However, this is not true, as we see below in Example 6.7. This fact motivates the need for Algorithm 1 in the next subsection.
Example 6.7 (Zero-divisor ACR is not detected through Gröbner bases).
Consider the following polynomials:
$f_{1}~{}$ $=~{}x_{B}x_{C}[(x_{A}-1)^{2}(x_{B}+3)+(x_{C}-2)]$
$f_{2}~{}$ $=~{}x_{B}x_{C}[(x_{A}-1)(x_{B}+3)^{2}-(x_{C}-2)]~{}.$
These equations lead to a mass-action system in three species $A$, $B$, and $C$ with $dx_{A}/dt=f_{1}$, $dx_{B}/dt=dx_{C}/dt=f_{2}$ (recall Remark 2.10). The reduced Gröbner basis of the steady-state ideal $I=\langle f_{1},f_{2}\rangle$ with respect to the lexicographic order $x_{B}>x_{C}>x_{A}$ (which is an elimination order for $\{x_{B},x_{C}\}$, that is, for eliminating $x_{A}$) is ${\mathcal{G}}=\{g_{1},g_{2},g_{3},g_{4}\}$, where
(25) $g_{1}$ $=x_{B}x_{C}(x_{C}-2)(-x_{C}+x_{A}^{3}-3x_{A}^{2}+3x_{A}+1),$
$g_{2}$ $=x_{B}x_{C}(x_{B}x_{A}^{2}-2x_{B}x_{A}+x_{B}+x_{C}+3x_{A}^{2}-6x_{A}+1),$
$g_{3}$ $=x_{B}x_{C}(-x_{B}x_{C}+3x_{B}x_{A}^{2}-6x_{B}x_{A}+5x_{B}-x_{C}x_{A}+x_{C}+9x_{A}^{2}-16x_{A}+7),$
$g_{4}$ $=x_{B}x_{C}(x_{B}^{2}x_{A}-x_{B}^{2}+6x_{B}x_{A}-6x_{B}-x_{C}+9x_{A}-7).$
Observe that ${\mathcal{G}}\cap{\mathbb{Q}}[x_{A}]=\varnothing$ (and so $x_{A}-1$ is not in $I$) and also that $x_{A}-1$ does not divide any element of ${\mathcal{G}}$.
To confirm that this example shows that condition (24) is not necessary for zero-divisor ACR, it suffices to show that there is zero-divisor ACR in $A$ with ACR-value $1$. To see that $x_{A}-1$ is a zero-divisor of $I$, observe that the following polynomial belongs to $I=\langle f_{1},f_{2}\rangle$:
$h~{}=~{}f_{1}+f_{2}~{}=~{}(x_{A}-1)x_{B}x_{C}(x_{B}+3)[(x_{A}-1)+(x_{B}+3)]~{}.$
We saw earlier that $x_{A}-1$ is not in $I$, and it is straightforward to check that $h/(x_{A}-1)=x_{B}x_{C}(x_{B}+3)[(x_{A}-1)+(x_{B}+3)]$ also is not in $I$. Hence, $x_{A}-1$ is a zero-divisor of $I$.
Next, from examining the factors of $h$, we see that every positive steady state satisfies $x_{A}=1$. We use this fact, together with the structure of the polynomials $f_{1}$ and $f_{2}$, to see that the positive steady-state locus is $\{(1,x_{B},2)\mid x_{B}>0\}$. This confirms that there is zero-divisor ACR in species $A$ with ACR-value $1$ (and also ACR in $C$).
6.2. Algorithm
The ideas in the prior subsection motivate Algorithm 1, which outputs candidates for ACR species (and the corresponding ACR-values).
Input: A mass-action system $(G,\kappa^{*})$ with $n$ species (and rate constants $\kappa^{*}$ in $\mathbb{Q}$).
Output: A finite set $S$ of pairs $(X_{j},\beta)$.
Initialize $S=\varnothing$;
Let $I$ be the steady-state ideal of $(G,\kappa^{*})$;
for $i\in\{1,2,\dots,n\}$ do
       Call $\hat{\mathbf{x}}_{i}=(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n})$. Let $\mathcal{G}$ be a reduced Gröbner basis of $I$ for an elimination order for $\hat{\mathbf{x}}_{i}$ on ${\mathbb{Q}}[x_{1},x_{2},\dots,x_{n}]$. Let ${\mathcal{G}}_{i}:={\mathcal{G}}\cap{\mathbb{Q}}[x_{i}]$;
       if ${\mathcal{G}}_{i}\neq\varnothing$ then
            Let $\widetilde{h}$ denote the unique element of ${\mathcal{G}}_{i}$;
             for $\alpha$ a positive root of $-~\widetilde{h}$ do
                  append $(X_{i},\alpha)$ to $S$
      else
            for $g\in\mathcal{G}$ do
                   for $\alpha$ a positive root of the leading coefficient ${\rm lc}_{\hat{\mathbf{x}}_{i}}(g)$ do
                        append $(X_{i},\alpha)$ to $S$
Algorithm 1 Candidates for ACR
We conjecture, as follows, that Algorithm 1 succeeds for all systems with zero-divisor ACR.
Conjecture 6.8.
Consider a mass-action system $(G,\kappa^{*})$ with rational-number rate constants. If $(G,\kappa^{*})$ has zero-divisor ACR in species $X_{i}$ with ACR-value $\alpha$, then $(X_{i},\alpha)$ is one of the outputs of Algorithm 1.
Remark 6.9.
Recall from Definition 6.2 that zero-divisor ACR means that either (1) $x_{i}-\alpha$ is in the steady-state ideal or (2) $x_{i}-\alpha$ is a zero-divisor of the steady-state ideal. In case (1), it is straightforward to see that Conjecture 6.8 holds. Indeed, in that case, $x_{i}-\alpha$ is the unique element of ${\mathcal{G}}\cap{\mathbb{Q}}[x_{i}]$, and so $(X_{i},\alpha)$ is an output of the algorithm.
Next, we present an example that shows how to apply Algorithm 1 and interpret its output.
Example 6.10 (Example 6.1, continued).
Recall that the network from Example 6.1 has steady-state ideal $I=\langle-x_{A}x_{C}x_{D}+x_{C}x_{D},~{}x_{C}-x_{B},~{}x_{A}^{2}x_{C}-3x_{A}x_{C}+2x_{C},~{}x_{A}^{2}x_{D}-4x_{A}x_{D}+3x_{D}\rangle$.
  • The reduced Gröbner basis $\mathcal{G}$ of $I$ with respect to the lexicographic order $x_{B}>x_{C}>x_{D}>x_{A}$ was shown in (23), where we obtained ${\mathcal{G}}\cap{\mathbb{Q}}[x_{A}]=\varnothing$. Now Algorithm 1 asks us to find the positive roots of the leading coefficients with respect to $\mathbf{x}=(x_{B},x_{C},x_{D})$; these coefficients are $x_{A}^{2}-4x_{A}+3,x_{A}^{2}-3x_{A}+2,x_{A}-1,1$. Therefore, the algorithm outputs three pairs: $(x_{A},1)$, $(x_{A},3)$, and $(x_{A},2)$.
  • If we consider the lexicographic order $x_{A}>x_{C}>x_{D}>x_{B}$, the reduced Gröbner basis of $I$ with respect to this order is
    ${\mathcal{G}}=\{x_{C}-x_{B},~{}x_{A}x_{B}x_{D}-x_{C}x_{D},~{}x_{A}^{2}x_{B}-3x_{A}x_{C}+2x_{C},~{}x_{A}^{2}x_{D}-4x_{A}x_{D}+3x_{D}\}.$
    As ${\mathcal{G}}\cap{\mathbb{Q}}[x_{B}]=\varnothing$, the algorithm asks us to find the positive roots of the leading coefficients with respect to $\mathbf{x}=(x_{A},x_{C},x_{D})$: $1$ and $x_{B}$. As there are no positive roots, the algorithm returns no candidates for ACR-values in $x_{B}$.
  • The situation for $x_{C}$ and $x_{D}$ is similar to $x_{B}$: no candidate ACR-values are returned.
Thus, for this system, Algorithm 1 returns three possible ACR-values for species $A$, namely, $1$, $2$, and $3$. Next, we show that, in fact, there is ACR in $A$, and the ACR-value is one of the candidates, namely, $1$. Indeed, it is straightforward to see that the positive steady-state locus is $\{(1,\alpha,\alpha,\beta)\mid\alpha>0,~{}\beta>0\}$. As we saw in Example 6.1, $x_{A}-1$ is a zero divisor of $I$ and hence the system has zero-divisor ACR in the species $A$.
In the prior example, as we perform the algorithm, we encounter the empty set for ${\mathcal{G}}_{i}:={\mathcal{G}}\cap{\mathbb{Q}}[x_{i}]$. In contrast, this set is nonempty in the next example.
Example 6.11.
[…]
The next example features a network with vacuous ACR.
Example 6.12.
[…]
Example 6.13 (Example 6.3, continued).
We return to the Shinar and Feinberg network (18). Symbolic rate constants appear in what follows, even though the algorithm applies to specialized values of rate constants, because the steps of the algorithm (in this case) do not depend on the specific choice of the rate constants. We apply the algorithm, as follows. Recall that the reduced Gröbner basis ${\mathcal{G}}$ of the steady-state ideal $I$ with respect to the lexicographic order $x_{1}>x_{2}>x_{4}>x_{6}>x_{7}>x_{5}$ satisfies ${\mathcal{G}}\cap{\mathbb{Q}}[x_{5}]=\varnothing$. We therefore consider the positive roots of the leading coefficients with respect to $\mathbf{x}=(x_{1},x_{2},x_{3},x_{4},x_{6},x_{7})$, which are $\kappa_{7}\kappa_{9}x_{5}-(\kappa_{3}\kappa_{8}+\kappa_{3}\kappa_{9}),\kappa_{6},\kappa_{4},\kappa_{3},\kappa_{1}$. We obtain $(X_{5},\frac{\kappa_{3}\kappa_{8}+\kappa_{3}\kappa_{9}}{\kappa_{7}\kappa_{9}})$ as an output of Algorithm 1. As we saw earlier, indeed $X_{5}=Y_{p}$ exhibits (zero-divisor) ACR.
Next, we return to an example in which zero-divisor ACR was not detected through the Gröbner bases of the steady-state ideal (Example 6.7). Fortunately, however, ACR detected through Algorithm 1.
Example 6.14 (Example 6.7, continued).
We apply Algorithm 1 to the mass-action system in Example 6.7. The leading coefficients of the Gröbner basis elements (25) with respect to $\mathbf{x}=(x_{B},x_{C})$ are $1$, $(x_{A}-1)^{2}$, $1$, and $x_{A}-1$, respectively. The algorithm therefore outputs the pair $(x_{A},1)$. (Recall that, indeed, there is ACR in species $A$.)
Next, we consider species $C$. In the reduced Gröbner basis of $I$ with respect to the lexicographic order $x_{A}>x_{B}>x_{C}$, the leading coefficients with respect to $\mathbf{x}=(x_{A},x_{B})$ are $x_{C}(x_{C}-2)$ and $x_{C}$. Hence, the algorithm returns the pair $(x_{C},2)$, which is the actual ACR-value for species $C$, as we saw in Example 6.7. In fact, $x_{C}-2$ is also a zero-divisor of $I$, which is straightforward to check from the first element of the Gröbner basis which is $x_{B}x_{C}(x_{C}-2)(x_{B}^{3}+9x_{B}^{2}+27x_{B}+x_{C}+25)$.
6.3. Connection to parametric ideals and Gröbner covers
In the future, we would like to resolve Conjecture 6.8. One approach to doing so is through the theory of Gröbner bases for ideals involving parameters, which in fact inspired Algorithm 1 in the first place. We describe this theory, as it pertains to the algorithm, in this subsection.
In general, a parametric ideal is an ideal $I\subseteq\mathbb{F}(t_{1},t_{2},\dots,t_{m})[x_{1},x_{2},\dots,x_{n}]$, where $t_{1},t_{2},\dots,t_{m}$ are the parameters. Given a Gröbner basis $\mathcal{G}$ of $I$, a fundamental problem is to find a proper variety $W\subseteq\mathbb{F}^{m}$ such that $\mathcal{G}$ remains a Gröbner basis under all specializations $(t_{1},t_{2},\dots,t_{m})\longmapsto(a_{1},a_{2},\dots,a_{m})\in\mathbb{F}^{m}\smallsetminus W$. Much theory has been developed to tackle this problem (see, for instance, the book [54]). A common approach to solve this problem involves turning the parameters into variables, that is, considering the ring $\mathbb{F}[\mathbf{x},\mathbf{t}]=\mathbb{F}[x_{1},x_{2},\dots,x_{n},t_{1},t_{2},\dots,t_{m}]$.
In this setting, a result of Suzuki and Sato is fundamental [73] (see also [54, Lemma 3.11] and [16, §6.3 Proposition 1]). We point out that their result holds for an arbitrary (finite) number of parameters $t_{1},t_{2},\dots,t_{m}$. However, to simplify the exposition, we restrict this result to the one-parameter case and rephrase it in our particular context which is when this parameter is in fact one of the variables $x_{i}$. The resulting Lemma 6.15 inspired our Algorithm 1.
Lemma 6.15 (Suzuki-Sato).
Given an ideal $I\subseteq{\mathbb{Q}}[x_{1},x_{2},\dots,x_{n}]$, fix $i\in\{1,2,\dots,n\}$ and call $\hat{\mathbf{x}}_{i}=(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n})$. Consider an elimination order $\succ$ for $\hat{\mathbf{x}}_{i}$ on ${\mathbb{Q}}[x_{1},x_{2},\dots,x_{n}]$, and let $\mathcal{G}=\{g_{1},g_{2},\dots,g_{s}\}$ be a reduced Gröbner basis of $I$ with respect to this order. If $\alpha\in\overline{{\mathbb{Q}}}$ is such that $lc_{\hat{\mathbf{x}}_{i}}(g_{j})|_{x_{i}=\alpha}\neq 0$ for any $1\leq j\leq s$ with $g_{j}\notin\mathcal{G}\cap{\mathbb{Q}}[x_{i}]$, then the specialized set $\{g_{1}|_{x_{i}=\alpha},g_{2}|_{x_{i}=\alpha},\dots,g_{t}|_{x_{i}=\alpha}\}$ is a reduced Gröbner basis of the specialized ideal $I|_{x_{i}=\alpha}\subseteq\overline{{\mathbb{Q}}}[\hat{\mathbf{x}}_{i}]$ with respect to the order $\succ_{\hat{\mathbf{x}}_{i}}$.
Remark 6.16.
Lemma 6.15 extends to any computable field $\mathbb{F}$, in which case we consider $\alpha$ in the algebraic closure of $\mathbb{F}$ [54].
As explained above, Lemma 6.15 is part of the theory of parametric Gröbner bases. There are several algorithms for computing with parametric Gröbner bases, one of which is implemented in Singular [18] and is based on the theory of Gröbner covers [54, 55]. We end this section by illustrating this theory through the following example.
Example 6.17 (Example 6.14 continued; zero-divisor ACR detected through Gröbner covers).
We revisit the mass-action system that was first shown in Example 6.7, in which the steady-state ideal is generated by the following polynomials:
$f_{1}~{}$ $=~{}x_{B}x_{C}[(x_{A}-1)^{2}(x_{B}+3)+(x_{C}-2)]$
$f_{2}~{}$ $=~{}x_{B}x_{C}[(x_{A}-1)(x_{B}+3)^{2}-(x_{C}-2)]~{}.$
The following Singular code computes the Gröbner cover of $I=\langle f_{1},f_{2}\rangle$:
LIB "grobcov.lib";
ring r = (0,xA),(xB,xC),dp;
ideal I = xB*xC*((xA-1)^2*(xB+3)+(xC-2)), xB*xC*((xA-1)*(xB+3)^2+(xC-2));
grobcov(I,("rep",1));
[…]
  • “initial” is a monomial ideal which is the initial ideal of any specialization at any point in “segment”, and
  • “basis” is the corresponding reduced Gröbner basis.
The option (“rep”,1) returns the segments in canonical C-representation, that is, as pairs of the form $(J_{1},J_{2})$ representing the interval $V(J_{1})\smallsetminus V(J_{2})$, where the ideals $J_{1}$ and $J_{2}$ are radical and $J_{2}\supseteq J_{1}$.
In our case, the output shows that any specialization of the ideal $I$ at any point in the segment $V(0)\smallsetminus V(x_{A}-1)={\mathbb{Q}}\smallsetminus\{1\}$ has the initial ideal $\langle x_{B}^{2}x_{C},\,x_{B}x_{C}^{3}\rangle$. However, when $x_{A}=1$, the initial ideal is $\langle x_{B}x_{C}^{2}\rangle$. We recall from Example 6.7 that $x_{A}=1$ is precisely the ACR-value for the corresponding mass-action system. Thus, in this example, the presence of ACR is detected via Gröbner covers.
Example 6.17 motivates the following question: If a mass-action system has zero-divisor ACR, is this always detected via a Gröbner cover? We believe that an affirmative answer would be a key step toward resolving Conjecture 6.8, and therefore validating Algorithm 1.
7. Discussion
This paper gives a comprehensive overview into the algebra and geometry of absolute concentration robustness. While at first glance, the property of ACR seems like it should be straightforward to check algorithmically, this paper highlights the subtleties involved. Indeed, the problem of detecting ACR ends up to be a nuanced problem in real algebraic geometry: Does every positive real component of a complex algebraic variety lie in a hyperplane of the form $\{x_{i}-\alpha$}?
The motivation for this study is biochemical reaction networks; however, there may be other applications that distill down to the same question. Thus, this manuscript can serve as guide in those cases. For example, there are several techniques that would be good first steps, while others would be good second steps. Other techniques would be valuable for in-depth case studies of parameterized polynomial systems.
From the point-of-view of computational algebraic geometry, given a reaction system $(G,\kappa)$ with steady-state ideal $I$, the first step in checking for ACR would be through previously suggested methods as captured in Proposition 4.16 and Remark 6.6. In particular, we suggest to first compute a Gröbner basis of $I$ and check for an element of the form $x_{i}-\alpha$. In most examples, this was enough for us to detect ACR. If that doesn’t work, then we suggest computing the saturation $I:(x_{1}x_{2}\cdots x_{n})^{\infty}$ and checking for an element of the form $x_{i}-\alpha$. Then, if that doesn’t work, compute each elimination ideal $I\cap\mathbb{Q}[x_{i}]$ and check whether there is a generator $g$ of any of the elimination ideals such that $g$ has a unique positive root. While these three computations constitute reasonable first steps, as we saw through the Examples 4.184.20, none of the conditions in Proposition 4.16 are necessary conditions for ACR. Indeed, this point is the main motivation for this paper.
If the methods suggested by Proposition 4.16 fail, either due to inconclusive results or computational limitations, the next steps would be those proposed in Sections 4.36.2. If the system is small, then a next step could be Procedure 4.14, which relies on computing a decomposition of the positive-restriction ideal and successively augmenting the ideals with Jacobian minors. However, it may be that the system is too large to effectively compute decompositions in a reasonable time. In this case, or in the case that we have irrational rate constants $\kappa$ or we expect irrational ACR-value $\alpha$, then we suggest numerically checking for ACR through a numerical irreducible decomposition (Procedure 5.7). For those looking for a more complete algebraic or geometric description of a specific network in respect to ACR, then we suggest Algorithm 1, which examines the leading coefficients of generators of elimination ideals obtained through Gröbner bases, as well as Procedure 5.7, which returns a dense sample of each real component intersected with the positive orthant.
While we envision the techniques described in this paper being used more generally, there are some techniques that are uniquely suited to chemical reaction networks due to the wealth of existing literature. For chemical reaction networks, we suggest starting with the techniques in Section 3 before going on to analysis using computational algebraic geometry. For example, one-species networks are fully classified in [52]. For chemical reaction networks with more than one species, then we suggest using Proposition 3.3 first. In particular, if it is known a priori that the network is not multistationary and the stochiometric subspace is $\mathbb{R}^{n}$, where $n$ is the number of species, then ACR follows. After that, we suggest using Proposition 3.9. In this case, there are several network-based results that can be used to establish the existence of particular types of positive steady-state parameterizations, such as monomial and rational parameterizations, and to then construct them [36, 53, 56]. While Section 3 is based on known results, we are hoping that collecting them together before getting into the results of later sections helps those looking for a complete guide to the state-of-the-art regarding ACR.
Finally, Section 6 raises some interesting questions related to computational algebraic geometry and ACR. In particular, we asked whether zero-divisor ACR can always be detected from a Gröbner basis with respect to an elimination order (Conjecture 6.8). An affirmative answer would certify that Algorithm 1 always succeeds in detecting zero-divisor ACR, thereby strengthening the connection between ACR and computational algebraic geometry.
Acknowledgements
[…]
Appendix A The Real Nullstellensatz for rational coefficients
The Real Nullstellensatz (Proposition 2.28) is well known, but we did not readily find the version we need in the literature. Accordingly, the aim of this appendix is to show how Proposition 2.28 follows from a result in [50].
The authors of [50] considered an ordered field $\mathbf{K}$ and a real closed extension of $\mathbf{K}$, denoted by $\mathbf{R}$ (for instance, $\mathbf{K}=\mathbb{Q}$ and $\mathbf{R}=\mathbb{R}$). Next, they defined the following subset of $\mathbf{K}[x_{1},x_{2},\dots,x_{n}]$:
$\mathcal{N}(\varnothing)~{}:=~{}\left\{\sum_{i=1}^{m}\omega_{i}V_{i}^{2}~{}|~{}m\in\mathbb{Z}_{\geq 0},~{}\omega_{1},\omega_{2},\dots,\omega_{m}\in\mathbf{K}_{>0},\,V_{1},V_{2},\dots,V_{m}\in\mathbf{K}[x_{1},x_{2},\dots,x_{n}]\right\}~{}.$
With this notation, the following result is called the Real Nullstellensatz in [50]:
Proposition A.1 (Theorem 1.11 in [50]).
Let $Q,P_{1},P_{2},\dots,P_{s}\subseteq\mathbf{K}[x_{1},x_{2},\dots,x_{n}]$, and let $J$ denote the ideal in $\mathbf{K}[x_{1},x_{2},\dots,x_{n}]$ generated by $P_{1},P_{2},\dots,P_{s}$. If $Q$ vanishes on the common zero set of $P_{1},P_{2},\dots,P_{s}$ in $\mathbf{R}^{n}$, then there exist $e\in\mathbb{Z}_{>0}$, $N\in\mathcal{N}(\varnothing)$, and $Z\in J$ such that
$Q^{2e}+N=Z~{}.$
Now we use Proposition A.1 (specifically, the case of $\mathbf{K}=\mathbb{Q}$) to prove Proposition 2.28.
Proof of Proposition 2.28.
Let $J$ be an ideal of $\mathbb{Q}[x_{1},x_{2},\dots,x_{n}]$. We must show that $\mathbb{R}]{J}$ (the real radical of $J$) equals the vanishing ideal of $V_{\mathbb{R}}(J)$ (the real variety of $J$).
For the first containment, let $P\in\sqrt[\mathbb{R}]{J}$ and let $x\in V_{\mathbb{R}}(J)$. We must show that $P(x)=0$. As $P\in\sqrt[\mathbb{R}]{J}$, there exist $m,\ell\in\mathbb{Z}_{\geq 0}$ and $h_{1},h_{2},\dots,h_{\ell}\in{\mathbb{Q}}[x_{1},x_{2},\dots,x_{n}]$ such that $P^{2m}+\sum_{i=1}^{\ell}h_{i}^{2}\in J$. Thus, as $x\in V_{\mathbb{R}}(J)$, we have the following sum of squares:
$(P(x))^{2m}+\sum_{i=1}^{\ell}(h_{i}(x))^{2}~{}=~{}0~{}.$
We conclude, as desired, that $P(x)=0$.
For the reverse containment, consider the ordered field $\mathbf{K}={\mathbb{Q}}$. Every positive rational number is a sum of squares, as follows:
$\frac{p}{q}~{}=~{}p\cdot q\cdot\frac{1}{q^{2}}~{}=~{}\underset{p\cdot q\text{ times}}{\underbrace{\frac{1}{q^{2}}+\dots+\frac{1}{q^{2}}}}~{},$
for $p,q\in{\mathbb{Z}}_{>0}$. This fact readily implies that $\mathcal{N}(\varnothing)$ consists of sums of squares, as follows: $\mathcal{N}(\varnothing)=\{\sum_{i=1}^{m}V_{i}^{2}~{}|~{}V_{i}\in{\mathbb{Q}}[x_{1},x_{2},\dots,x_{n}]\}$. This property of $\mathcal{N}(\varnothing)$, together with Proposition A.1 (applied to $\mathbf{K}=\mathbb{Q}$ and $\mathbf{R}=\mathbb{R}$), imply the following: for every $P$ in the vanishing ideal of $V_{\mathbb{R}}(J)$, there exist $e,m\in\mathbb{Z}_{>0}$ and $V_{1},V_{2},\dots,V_{m}\in{\mathbb{Q}}[x_{1},x_{2},\dots,x_{n}]$ such that the sum $P^{2e}+\sum_{i=1}^{m}V_{i}^{2}$ is in $J$, that is, $P\in\sqrt[\mathbb{R}]{J}$. Hence, the ideal of $V_{\mathbb{R}}(J)$ is contained in $\mathbb{R}]{J}$. ∎
Appendix B Complex-number ACR
In this appendix, we widen our scope to the (algebraically closed field of the) complex numbers ${\mathbb{C}}$, where in general we lose chemical significance, but gain much from the theoretical and computational point of view. Specifically, we introduce the notion of complex-number absolute concentration robustness (CACR), which – in contrast to ACR – can be readily detected from the steady-state ideal (Proposition B.6 and Algorithm 2).
In short, CACR generalizes the notion of ACR, by broadening the consideration from all positive zeros to all complex zeros with nonzero coordinates.
Definition B.1 (CACR for mass-action systems).
A mass-action system $(G,\kappa)$ with $n$ species and mass-action ODEs as in (2), has:
  1. (1)
    complex-number absolute concentration robustness (CACR) in species $X_{i}$ if, for every $x^{*}\in({\mathbb{C}}^{*})^{n}$ satisfying $f_{\kappa}(x^{*})_{1}=f_{\kappa}(x^{*})_{2}=\dots=f_{\kappa}(x^{*})_{n}=0$, the value of $x_{i}$ is the same. This value of $x_{i}$ is the CACR-value.
  2. (2)
    vacuous CACR if there does not exist $x^{*}\in({\mathbb{C}}^{*})^{n}$ satisfying $f_{\kappa}(x^{*})_{1}=f_{\kappa}(x^{*})_{2}=\dots=f_{\kappa}(x^{*})_{n}=0$.
As we do for ACR, we say that $(G,\kappa)$ has CACR if it has CACR in some species. It is immediate from the relevant definitions that CACR implies (possibly vacuous) ACR.
Proposition B.2 (CACR implies ACR).
If a mass-action system $(G,\kappa)$ has CACR in some species $X_{i}$, then the system also has ACR in $X_{i}$. Moreover:
  1. (1)
    If $(G,\kappa)$ has non-vacuous CACR in $X_{i}$ with CACR-value a positive real number, then $(G,\kappa)$ has (possibly vacuous) ACR in $X_{i}$.
  2. (2)
    If $(G,\kappa)$ has vacuous CACR or has non-vacuous CACR in $X_{i}$ with CACR-value that is not a positive real number, then $(G,\kappa)$ has vacuous ACR.
Example B.3 (Example 3.6, continued).
Recall that every mass-action system generated by the following network has ACR in species $A$ (with ACR-value $a^{*}=\kappa_{2}/\kappa_{1}$):
$G~{}=~{}\left\{A+B\overset{\kappa_{1}}{\underset{\kappa_{2}}{\leftrightarrows}}B\overset{\kappa_{3}}{\underset{\kappa_{4}}{\leftrightarrows}}0~{},~{}2B\overset{\kappa_{5}}{\underset{\kappa_{6}}{\leftrightarrows}}3B\right\}~{}.$
It is straightforward to check that these systems also have CACR in species $A$ (with CACR-value $x_{A}^{*}=\kappa_{2}/\kappa_{1}$).
Example B.4 (Example 4.15, continued).
[…]
Another example with ACR but not CACR is Example 2.9.
What we gain from expanding the definition of ACR to CACR is necessary and sufficient conditions for CACR. Moreover, in theory, CACR can be detected algorithmically (if computations over $\mathbb{C}$ are possible). We record this claim in Proposition B.6 and Algorithm 2 below, which make use of radical ideals (7) and saturation ideals (Definition 2.22). We also extend steady-state ideals to reside over $\mathbb{C}$ as follows:
Definition B.5.
For a mass-action system $(G,\kappa)$, the complex-steady-state ideal is the ideal of $\mathbb{C}[x_{1},x_{2},\ldots,x_{n}]$ generated by the right-hand sides of the ODEs in (2).
For part (3) of the next result, recall that any ideal of $\mathbb{C}[x_{i}]$ (with only one variable $x_{i}$) is principally generated, so we may speak of “a generator”. Also, the proof makes use of Hilbert’s Nullstellensatz (e.g., see [16]).
Proposition B.6 (CACR and ideals).
Let $(G,\kappa^{*})$ be a mass-action system with $n$ species and complex-steady-state ideal $I$. Let $\mathfrak{m}:=x_{1}x_{2}\cdots x_{n}$, and let $1\leq i\leq n$. Then $(G,\kappa^{*})$ has vacuous CACR if and only if $\sqrt{(I:\mathfrak{m}^{\infty})}=\langle 1\rangle$. On the other hand, for $\alpha\in\mathbb{C}^{*}$, the following are equivalent:
  1. (1)
    $(G,\kappa^{*})$ has non-vacuous CACR in species $X_{i}$ with CACR-value $\alpha$,
  2. (2)
    $x_{i}-\alpha$ is in $\sqrt{(I:\mathfrak{m}^{\infty})}$, and $\sqrt{(I:\mathfrak{m}^{\infty})}\neq\langle 1\rangle$,
  3. (3)
    $x_{i}-\alpha$ is a generator of $\sqrt{(I:\mathfrak{m}^{\infty})\cap\mathbb{C}[x_{i}]}$.
Proof.
The equivalence of vacuous CACR and $\sqrt{(I:\mathfrak{m}^{\infty})}=\langle 1\rangle$ is readily deduced from [16, Theorem 10 of § 4] and the Weak Nullstellensatz. It is also straightforward to see that (2) and (3) are equivalent, so we show the equivalence of (1) and (2) below.
First assume that the system has non-vacuous CACR in species $X_{i}$ with CACR-value $\alpha$. Then, $\mathfrak{m}\,(x_{i}-\alpha)$ is identically zero over all the zeros of $I$ in ${\mathbb{C}}^{n}$. We deduce from Hilbert’s Nullstellensatz that there exists a natural number $k$ such that $\mathfrak{m}^{k}\,(x_{i}-\alpha)^{k}\in I$. Thus, $(x_{i}-\alpha)^{k}\in(I:\mathfrak{m}^{\infty})$ and so $x_{i}-\alpha\in\sqrt{(I:\mathfrak{m}^{\infty})}$.
Now assume that $x_{i}-\alpha$ is in $\sqrt{(I:\mathfrak{m}^{\infty})}$. Then $\mathfrak{m}^{k}(x_{i}-\alpha)^{k}\in I$ for some $k$. Consider $\tilde{x}^{*}\in({\mathbb{C}}^{*})^{n}$ for which $f_{\kappa}(\tilde{x}^{*})_{1}=f_{\kappa}(\tilde{x}^{*})_{2}=\dots=f_{\kappa}(\tilde{x}^{*})_{n}=0$. It follows that $(\tilde{x}^{*}_{1}\tilde{x}^{*}_{2}\dots\tilde{x}^{*}_{n})^{k}(\tilde{x}^{*}_{i}-\alpha)^{k}=0$. As $(\tilde{x}^{*}_{1}\tilde{x}^{*}_{2}\dots\tilde{x}^{*}_{n})\neq 0$, we have $(\tilde{x}^{*}_{i}-\alpha)^{k}=0$, and so $\tilde{x}^{*}_{i}=\alpha$. Thus, we have CACR in $X_{i}$ with CACR-value $\alpha$. ∎
Proposition B.6 yields Algorithm 2 for detecting CACR. The algorithm does not compute the radical ideal until Step 3 because it is cheaper to do computations in one variable. For details about the computations of these ideals and generators, see the book [16].
Input: The complex-steady-state ideal $I$ of a mass-action system $(G,\kappa^{*})$. Output: The CACR-value if $(G,\kappa^{*})$ has non-vacuous CACR in $X_{i}$; “Vacuous ACR” or “No CACR” if $(G,\kappa^{*})$ has, respectively, vacuous ACR or no CACR in $X_{i}$. Step 0: Let $\mathfrak{m}:=x_{1}x_{2}\dots x_{n}$, where $n$ is the number of species of $G$. Step 1: Compute the saturation ideal $(I:\mathfrak{m}^{\infty})$. Step 2: Compute the elimination ideal $I_{i}:=(I:\mathfrak{m}^{\infty})\cap\mathbb{C}[x_{i}]$. Step 3: Compute $\sqrt{I_{i}}$ and find a generator $g$. Step 4: If $g$ has degree 1, return its root. If $g$ has degree 0, return “Vacuous CACR”. Otherwise, return “No CACR”.
Algorithm 2 Algorithm for CACR.
Proof of correctness of Algorithm 2.
The correctness of this algorithm follows directly from Proposition B.6 and the fact that the generator $g$ in Step 3 can not be the degree-1 polynomial $g=x_{i}$, due to having performed saturation in Step 1. ∎